### v1

# On stable assignments generated by choice functions of mixed type

Abstract
We consider one variant of stable assignment problems in a bipartite graph endowed with nonnegative capacities on the edges and quotas on the vertices. It can be viewed as a generalization of the stable allocation problem introduced by Bai̋ou and Balinsky, which arises when strong linear orders of preferences on the vertices in the latter are replaced by weak ones. At the same time, our stability problems can be stated in the framework of a theory by Alkan and Gale on stable schedule matchings generated by choice functions of a wide class. In our case, the choice functions are of a special, so-called mixed, type. The main content of this paper is devoted to a study of rotations in our mixed model, functions determining “elementary” transformations of one stable assignments into others; they look more sophisticated compared with rotations in the stable allocation problem that are generated by simple cycles. On this way, we explicitly construct a weighted poset of rotations and show that the stable assignments are bijective to the so-called closed functions for this poset.
Keywords: stable marriage, allocation, diversification, choice function, rotation
1 Introduction
Since the famous work of Gale and Shapley [7] on stable matchings (“marriages”) in bipartite graphs, there have appeared a variety of researches on models of stability on graphs. A well-known model among them was introduced by Bai̋ou and Balinsky [2] under the name of stable allocation problem.
In a setting of that problem, there is given a bipartite graph $G=(V,E)$ whose vertices are partitioned into two subsets (color classes), say, $F$ and $W$, interpreted as the sets of firms and workers. An edge $e\in E$ connecting vertices (or “agents”) $u$ and $v$ represents a possible “contract” between them, and its intensity (or an assignment on $e$) is restricted by a nonnegative real upper bound (or capacity) $b(e)\in{\mathbb{R}}_{+}$. Besides, each vertex $v\in V$ is endowed with a quota $q(v)\in{\mathbb{R}}_{+}$ which restricts the total assignment on the set $E_{v}$ of edges incident to $v$, and there is a (strong) linear order $>_{v}$ establishing preferences on the edges in $E_{v}$, namely, $e>_{v}e^{\prime}$ means that “agent” $v$ prefers “contract” $e$ to “contract” $e^{\prime}$. An assignment (or allocation) $x:E\to{\mathbb{R}}_{+}$ satisfying the capacity and quota constraints is called stable if, roughly, there is no edge $e=\{u,v\}$ such that $x(e)$ can be increased, without violation of the capacities and quotas anywhere, possibly decreasing $x$ on an edge $e^{\prime}\in E_{u}$ with $e^{\prime}<_{u}e$ and/or an edge $e^{\prime\prime}\in E_{v}$ with $e^{\prime\prime}<_{v}e$.
Bai̋ou and Balinsky [2] proved that a stable allocation always exists and can be found by an efficient algorithm. Also they showed that if the capacities and quotas are integer-valued, then there exists an integral stable allocation, and obtained some other impressive results. (Hereinafter an algorithm dealing with $G=(V,E)$ is said to be efficient if it runs in time polynomial in $|E|$, assuming w.l.o.g. that $|E|\geq|V|-1$.)
Alkan and Gale [1] considered much more general problem in a bipartite graph with capacities $b$ and, possibly, quotas $q$ as above. Here for each $v\in V$, there is a choice function $C_{v}$ that maps the set ${\cal B}_{v}:=\{z\in{\mathbb{R}}_{+}^{E_{v}}\colon z(e)\leq b(e)\;\forall e\in E_{v}\}$ into itself. The role of $C_{v}$ is twofold; in the case with quotas, this can be briefly explained as follows. First, when the amount $|z|:=\sum(z(e)\colon e\in E_{v})$ exceeds the quota $q(v)$, it reduces $z$ so as to get the equality $|z^{\prime}|=q(v)$, where $z^{\prime}=C_{v}(z)$. Second, it gives a way to compare the vectors in ${\cal B}_{v}$ obeying the quotas, by use of the preference relation $\succeq\,=\,\succeq_{v}$ defined by $z\succeq z^{\prime}\Longleftrightarrow C_{v}(z\vee z^{\prime})=z$ (where $z\vee z^{\prime}$ takes the values $\max\{z(e),z^{\prime}(e)\}$). By imposing to the choice functions $C_{v}$, $v\in V$, the axioms of consistence and persistence (which have been known, sometimes under different names, for simpler models, going back to Kelso and Crawford [13], Roth [14], Blair [3], and some others) and adding to them the axiom of size-monotonicity (cf. Fleiner [6]) or, stronger, of quota-filling, they succeeded to build a rather powerful and elegant theory of stability for weighted bipartite graphs. A central result in it (generalizing earlier results for particular cases) is that the set ${\cal S}$ of stable assignments (or stable schedule matchings, in terminology of [1]) is nonempty and represents a distributive lattice under the preference relations $\succeq_{F}$ determined by the choice functions $C_{v}$ with $v$ contained in the vertex part $F$ of $G$; namely, $x,y\in{\cal S}$ satisfy $x\succeq_{F}y$ if and only if for each $f\in F$, the restrictions $z,z^{\prime}$ of $x,y$ to $E_{v}$ (respectively) satisfy $z\succeq_{f}z^{\prime}$.
As an illustration in the setting with quotas, Alkan and Gale pointed out two special cases (Examples 1,2 in Sect. 2 of [1]). One of them is just the above-mentioned stable allocation problem, that we will briefly denote as SAP. In this case, for each $v\in V$, the choice function $C_{v}$ acts on $z\in{\cal B}_{v}$ with $|z|>q(v)$ so as to respect more preferred and ignore less preferred edges in $E_{v}$; more precisely, it selects the element $\widehat{e}\in E_{v}$ such that $p:=\sum(z(e)\colon e>_{v}\widehat{e})<q(v)$ and $p+z(\widehat{e})\geq q(v)$, preserves the values for all $e>_{v}\widehat{e}$, reduces to zero the values for all $e<_{v}\widehat{e}$, and changes the value on $\widehat{e}$ to $q(v)-p$. The second special case deals with the so-called diversifying choice function for each $v\in V$, which uniformly cuts values of $z\in{\cal B}_{v}$ with $|z|>q(v)$ from above; more precisely, it computes the number (“level”) $r$ such that $\sum(\min\{z(e),r\}\colon e\in E_{v})=q(v)$, and changes $z(e)$ to $\min\{z(e),r\}$, $e\in E_{v}$. In fact, this corresponds to the equivalence (or disorder) relation between all elements of $E_{v}$ from the viewpoint of the “agent” $v$, in contrast to the allocation case where $E_{v}$ is totally ordered. Surprisingly, the diversifying model is poor: for any $b$ and $q$, there exists exactly one stable solution, see [12].
In this paper, we consider a common generalization of SAP and the stable diversifying problem. It is obtained from SAP by admitting that for each $v\in V$, the preferences on $E_{v}$ are established by a weak linear order. Equivalently, the set $E_{v}$ is partitioned into subsets $\pi^{1},\ldots,\pi^{k}$, called ties, such that $e\in\pi^{i}$ prefers $e^{\prime}\in\pi^{j}$ if $i<j$, and $e,e^{\prime}$ are regarded as equivalent if $i=j$. We write $e>_{v}e^{\prime}$ if $i<j$, and $e\sim_{v}e^{\prime}$ if $i=j$. Then an assignment $x:E\to{\mathbb{R}}_{+}$ satisfying the capacities and quotas is stable if, roughly, there is no edge $e=\{u,v\}$ such that $x(e)$ can be increased, without violation of the capacities and quotas anywhere, possibly decreasing $x$ on an edge $e^{\prime}\in E_{u}$ such that either $e^{\prime}<_{u}e$, or $e^{\prime}\sim_{u}e$ and $x^{\prime}(e)>x(e)$, and/or an edge $e^{\prime\prime}\in E_{v}$ such that either $e^{\prime\prime}<_{v}e$, or $e^{\prime\prime}\sim_{v}e$ and $x^{\prime\prime}(e)>x(e)$. This case is within the framework of Alkan-Gale’s model as well; here for $v\in V$, the choice function $C_{v}$ respects the assignments on more preferred ties, ignore less preferred ties, and acts as in the diversifying case for one tie in between those (a precise definition will be given later). We refer to a $C_{v}$ of this sort as a choice function of mixed type, and denote the arising stable assignment problem as SMP.
We will take advantages from both ways to state SMP: via weak linear orders (a combinatorial setting), and via choice functions of mixed type (an analytic setting). On the second way, relying on Alkan-Gale’s theory, we utilize the fact that the set ${\cal S}={\cal S}^{\rm mix}(G,b,q,\geq)$ of stable assignments in SMP is nonempty and forms a distributive lattice under the corresponding preference relations on ${\cal S}$, appeal to the construction of the $F$-optimal stable assignment, that we denote as $x^{\rm min}$, and use some other results from [1].
The main content of this paper is devoted to a study of rotations, “elementary augmentations” connecting close stable assignments in SMP, which clarifies, to some extent, the structure of the lattice on ${\cal S}$. This is already beyond the content of [1], but is parallel, in a sense, to a study of rotations for SAP given by Dean and Munshi [4]. (Note that originally nice properties of rotations and their applications have been revealed in connection with the stable marriage and roommate problems in [8, 9, 10]. One can also mention a recent work [5] on the Boolean variant of Alkan-Gale’s model, where rotations and their posets are extensively studied.)
It turns out that the formation and properties of rotations in SMP are more intricate compared with SAP. In particular, while a rotation in SAP is generated by a simple cycle whose edges have unit weights, a rotation in SMP is represented by an integer-valued function $\rho$ on the edges of a certain strongly connected subgraph $L$, and it is possible that $\rho$ takes large values and $L$ has large vertex degrees. As a consequence, while SAP with integral $b,q$ always has an integral stable solution, SMP with integral $b,q$ does not guarantee the existence of a stable solution with small denominators; moreover, the convex hull of ${\cal S}$ may have vertices with exponentially large denominators. Nevertheless, we can explicitly construct, in time polynomial in $|E|$, the set of all possible rotations $\rho$ in SMP, along with their maximal possible weights $\tau(\rho)$ (that can be used in “elementary” transformations $x\rightsquigarrow x^{\prime}=x+\tau(\rho)\rho$ of stable assignments). Moreover, by an analogy with SAP, we can efficiently construct a poset of rotations $\rho$ endowed with weights $\tau(\rho)$ and show that the so-called closed functions associated with this poset one-to-one correspond to the stable assignments in SMP (thus obtaining a sort of “explicit representation” for ${\cal S}$).
This paper is organized as follows. Section 2 contains basic definitions and settings and reviews important results from [1] needed to us. Here we recall that in a general case: the set ${\cal S}$ of stable assignments is nonempty; the partial orders $\succeq_{F}$ and $\succeq_{W}$ on ${\cal S}$ that express preferences for the parts $F$ (“firms”) and $W$ (“workers”) are polar to each other, i.e., for $x,y\in{\cal S}$, if $x\succeq_{F}y$ then $y\succeq_{W}x$, and vice versa; and $({\cal S},\succeq_{F})$ is a distributive lattice. In this lattice, we denote the best assignment for $F$ (and the worst for $W$) by $x^{\rm min}$, and the worst assignment for $F$ by $x^{\rm max}$. In Section 3 we recall the general method from [1] to construct a sequence of assignments that converges to $x^{\rm min}$. This method may take infinitely many iterations even if we deal with our model SMP. In the Appendix we devise an efficient modification to find $x^{\rm min}$ in SMP.
Section 4 is devoted to an in-depth study of rotations and related aspects in SMP. Here for the lattice $({\cal S},\succeq_{F})$ and an element $x\in{\cal S}-\{x^{\rm max}\}$, we consider the set ${\cal C}(x)$ of vectors (“elementary augmentations”) $\Delta\in{\mathbb{R}}^{E}$ such that $|\Delta|$ is sufficiently small and $x^{\prime}:=x+\Delta$ is a stable assignment satisfying $x\succ_{F}x^{\prime}$. The set ${\cal C}(x)$ is of conic type, in the sense that $\Delta\in{\cal C}(x)$ and $0\leq\varepsilon\leq 1$ imply $\varepsilon\Delta\in{\cal C}(x)$, and the irreducible vectors in ${\cal C}(x)$ can be characterized combinatorially. More precisely, we associate with $x$ a certain directed graph $\overrightarrow{\Gamma}(x)$ whose underlying undirected graph is a subgraph of $G$, which possesses the following property: each maximal strongly connected component of $\overrightarrow{\Gamma}(x)$ is bijective to an irreducible vector in ${\cal C}(x)$ considered up to proportionality. (Here a strongly connected, or, simply, strong component $L$ of $\overrightarrow{\Gamma}(x)$ is called maximal if no other strong component of $\overrightarrow{\Gamma}(x)$ is reachable by a directed path from $L$.) We denote by ${\cal L}^{\rm rot}(x)$ the set of maximal strong components in $\overrightarrow{\Gamma}(x)$, and associate with each $L=(V_{L},E_{L})\in{\cal L}^{\rm rot}$ a certain integer vector $\rho_{L}\in{\mathbb{Z}}^{E}$ having the support $E_{L}$ and representing a sort of integer circulation.
This vector $\rho_{L}$ is just what we mean under the rotation for $x$ generated by $L\in{\cal L}^{\rm rot}(x)$. One shows that the vectors $\varepsilon\rho_{L}$, $L\in{\cal L}^{\rm rot}(x)$, with a sufficiently small $\varepsilon>0$, are exactly the irreducible vectors in ${\cal C}(x)$ (cf. Corollary 4.7). (Note that in SAP each rotation $\rho$ is generated by a simple cycle and $\rho$ takes values $0,\pm 1$. In contrast, in SMP maximal strong components $L$ in $\overrightarrow{\Gamma}(x)$ may be rather cumbersome and $\rho_{L}$ may take big values; an example is illustrated in Sect. 5.) For $L\in{\cal L}^{\rm rot}(x)$, we define $\tau(\rho_{L})$ to be the maximal value $\lambda\geq 0$ such that $x+\lambda\rho_{L}$ is stable, and prove that for any $0<\lambda\leq\tau(\rho_{L})$, the assignment $x^{\prime}:=x+\lambda\rho_{L}$ is stable as well (cf. Lemma 4.8). We say that such an $x^{\prime}$ is obtained from $x$ by the rotational shift of weight $\lambda$ along $\rho_{L}$; when $\lambda=\tau(\rho_{L})$, the shift is called full. Another important fact shown in this section (generalizing a known analog for SAP) is that $x$ can be updated by using several rotations simultaneously; namely, for any $\lambda:{\cal L}^{\rm rot}(x)\to{\mathbb{R}}_{+}$ satisfying $\lambda(L)\leq\tau_{L}$ for all $L\in{\cal L}^{\rm rot}(x)$, the assignment $x^{\prime}:=x+\sum(\lambda(L)\rho_{L}\colon L\in{\cal L}^{\rm rot}(x))$ is stable, and $x\succeq_{F}x^{\prime}$ (cf. Proposition 4.9).
Section 5 discusses additional, including computational, aspects of rotations. We explain that for $x\in{\cal S}-\{x^{\rm max}\}$, the set of rotations $\rho=\rho_{L}$ applicable to $x$ (where $L\in{\cal L}^{\rm rot}$) and their maximal weights $\tau(\rho)$ can be found in time polynomial in $|E|$ (note that to construct the function $\rho_{L}$ from the component $L$, we have to solve a system of linear equalities). Also one shows that any $x\in{\cal S}$ can be reached from $x^{\rm min}$ by a sequence of $O(|E|)$ rotational shifts. The section finishes with an example (mentioned above) of rotation $\rho_{L}$ taking “big” values while $L$ is “small”.
[…]
In the Appendix, we elaborate an efficient algorithm to find the $F$-optimal stable assignment $x^{\rm min}$ in SMP. It modifies the method from Sect. 3, based on aggregation of “long” sequences of iterations in the latter, by solving certain linear programs.
2 Definitions, settings, backgrounds
$\bullet$  The stability problem SMP that we deal with is defined as follows. There is given a bipartite graph $G=(V,E)$ in which the set $V$ of vertices is partitioned into two subsets (color classes) $F$ and $W$, whose elements are conditionally called firms and workers, respectively. The edges $e\in E$ are equipped with upper bounds, or capacities, $b(e)\in{\mathbb{R}}_{+}$, and the vertices $v\in V$ with quotes $q(v)\in{\mathbb{R}}_{+}$. We will assume that all values $b(e)$ and $q(v)$ are positive; otherwise the edge/vertex can be removed without affecting the problem. One may interpret the vertices as “agents” in a market, and the edges as possible “contracts” between them whose values are bounded by the parameters $b$ and $q$. The set of edges incident to a vertex $v\in V$ is denoted by $E_{v}$.
For each vertex $v\in V$, there are preferences on the edges in $E_{v}$; they are described by a weak linear order, a transitive relation $\geq_{v}$ on pairs in $E_{v}$ in which different edges $e,e^{\prime}$ are admitted to satisfy simultaneously both comparisons $e\geq_{v}e^{\prime}$ and $e^{\prime}\geq_{v}e$. We write $e>_{v}e^{\prime}$ if $e\geq_{v}e^{\prime}$ but not $e^{\prime}\geq_{v}e$, which means that from the viewpoint of the vertex (“agent”) $v$, the edge $e$ is preferred to (or better than) the edge $e^{\prime}$; whereas if both $e\geq_{v}e^{\prime}$ and $e^{\prime}\geq_{v}e$ take place, then the edges $e,e^{\prime}$ are regarded by $v$ as equivalent, which is denoted as $e\sim_{v}e^{\prime}$.
It is useful for us to associate the weak order $\geq_{v}$ with partition $\Pi_{v}=(\pi^{1},\pi^{2},\ldots,\pi^{k})$ of the set $E_{v}$ such that $e\in\pi^{i}$ and $e^{\prime}\in\pi^{j}$ satisfy $e>_{v}e^{\prime}$ if $i<j$, and $e\sim_{v}e^{\prime}$ if $i=j$. We refer to $\Pi_{v}$ as a cortege, and to its members $\pi^{i}$ as ties, and may write $\pi^{1}>_{v}\pi^{2}>_{v}\cdots>_{v}\pi^{k}$. When $i<j$, we say that the tie $\pi^{i}$ is more preferable, or better, or earlier than $\pi^{j}$ (and $\pi^{j}$ is less preferable, or worse, or later than $\pi^{i}$).
$\bullet$  The set of admissible assignments for $(G,b)$ (without involving quotas $q$) is defined to be the box :=${\cal B}:=\{x\in{\mathbb{R}}_{+}^{E}\colon x\leq b\}$. For a vertex $v\in V$, the restriction of $x\in{\cal B}$ to the edge set $E_{v}$ is denoted by $x_{v}$; the set of these restrictions is denoted as ${\cal B}_{v}$. On this set, a choice function (CF) $C_{v}:{\cal B}_{v}\to{\cal B}_{v}$ acts as follows: for each $z\in{\cal B}_{v}$, $C_{v}(z)\leq z$, and $C_{v}$ is agreeable with the order $\geq_{v}$ and the quota $q(v)$, which means that:
  1. (2.1)
    • (a)
      if $|z|\leq q(v)$ then $C_{v}(z)=z$;
    • (b)
      if $|z|\geq q(v)$, then in the cortege $\Pi_{v}$, we take the tie $\pi^{i}$ such that $z(\pi^{1}\cup\cdots\cup\pi^{i-1})<q(v)\leq z(\pi^{1}\cup\cdots\cup\pi^{i})$, and define:
      $C_{v}(z)(e):=\begin{cases}z(e)&\mbox{if $e\in\pi^{j}$, $j<i$,}\\ 0&\mbox{if $e\in\pi^{j}$, $j>i$,}\\ r\wedge z(e)&\mbox{if $e\in\pi^{i}$;}\end{cases}$ (2.2)
      where $r=r^{z}$ (the cutting height in $\pi^{i}$) satisfies
      $z(\pi^{1}\cup\cdots\cup\pi^{i-1})+\sum(r\wedge z(e):e\in\pi^{i})=q(v).$
(Hereinafter, for real-valued functions $a,b$ on a set $S$, we use the standard notation $a\wedge b$ and $a\vee b$ for the functions taking values $\min\{a(e),b(e)\}$ and $\max\{a(e),b(e)\}$, $e\in S$, respectively. If $S^{\prime}$ is a finite subset of $S$, we write $a(S^{\prime})$ for the value $\sum(a(e):e\in S^{\prime})$. When $S$ is finite, $a(S)$ is also denoted as $|a|$.)
The tie $\pi^{i}$ figured in case ((2.1))(b) is called critical and denoted as $\pi^{\rm c}_{v}(z)$. The set of ties $\pi^{i}\in\Pi_{v}$ better (resp. worse) than $\pi^{\rm c}_{v}(z)$, i.e. $\pi^{i}>_{v}\pi^{\rm c}_{v}(z)$ (resp. $\pi^{i}<_{v}\pi^{\rm c}_{v}(z)$) is denoted by $\Pi^{\rm beg}_{v}(z)$ (resp. by $\Pi^{\rm end}_{v}(z)$). Note that $C_{v}(z)$ takes zero values on all edges of ties in $\Pi^{\rm end}_{v}(z)$.
We call $C_{v}$ defined as in ((2.1)) a choice function of the mixed type, or simply a mixed CF. Here the term “mixed” is justified by the fact that such a CF appears as a direct generalization of two cases known in the literature, pointed out as examples in [1].
Example 1. Let $\geq_{v}$ be a (strong) linear order on $E_{v}$; in other words, each tie in the cortege $\Pi_{v}$ consists of a unique element. Then the CFs $C_{v}$ as above are exactly those arising in the stable allocation problem of Bai̋ou and Balinsky [2] (denoted as SAP).
Example 2. If all edges in $E_{v}$ are equivalent (and therefore $\Pi_{v}$ has a unique tie), then $C_{v}$ is called a diversifying choice function, following terminology in [1].
$\bullet$  It is not difficult to check that a CF $C=C_{v}$ of mixed type satisfies the properties of consistence and persistence, in terminology of [1]; namely, for $z,z^{\prime}\in{\cal B}_{v}$, there hold:
  • (i)
    If $z\geq z^{\prime}\geq C(z)$, then $C(z^{\prime})=C(z)$ (consistence);
  • (ii)
    If $z\geq z^{\prime}$, then $C(z)\wedge z^{\prime}\leq C(z^{\prime})$ (persistence).
(One more axiom imposed in [1] is the size monotonicity one; its strengthening, the quota filling axiom, saying that $|z|\geq q(v)$ implies $|C(z)|=q(v)$, is automatically valid in our case.)
In particular, (i) and (ii) provide the equality $C(C(z))=C(z)$ for all $z\in{\cal B}_{v}$. A function (assignment) $z\in{\cal B}_{v}$ such that $z=C(z)$ is called rational; the set of these is denoted by ${\cal A}_{v}$. Following a general model of Alkan and Gale, the set ${\cal A}_{v}$ is endowed with the binary relation of “revealed preference” $\succeq$, where for $z,z^{\prime}\in{\cal A}_{v}$, we write $z\succeq z^{\prime}$ and say that $z$ is preferred to (or better than) $z^{\prime}$ if
$C(z\vee z^{\prime})=z.$
[…]
For further needs, it is useful to specify the features of relation $\succeq$ in our case. We say that $z\in{\cal A}_{v}$ is fully filled if $|z|=q(v)$, and deficit if $|z|<q(v)$. For a fully filled $z$, in the critical tie $\pi^{\rm c}_{v}(z)$ we distinguish the subset
$H_{v}(z):=\{e\in\pi^{\rm c}_{v}(z)\colon z(e)=r^{z}\},$ (2.3)
called the head of $z$. The set of edges occurring in the ties $\pi^{1},\pi^{2},\ldots,\pi^{\rm c}_{v}(x)$, except for those in $H_{v}(z)$, is denoted by $T_{v}(z)$ and called the tail of $z$. When $z$ is deficit, we formally define $H_{v}(z):=\emptyset$ and $T_{v}(z):=E_{v}$. One can check that
  1. (2.4)
    if both $z,z^{\prime}\in{\cal A}_{v}$ are fully filled, and $\pi^{i}$ is the critical tie $\pi^{\rm c}_{v}(z)$ for $z$, then $z\succeq z^{\prime}$ if and only if $z(e)\geq z^{\prime}(e)$ holds for all edges $e$ in the tail $T_{v}(z)$ (implying that $\pi^{\rm c}_{v}(z)\geq\pi^{\rm c}_{v}(z^{\prime})$, and if $\pi^{\rm c}_{v}(z^{\prime})=\pi^{\rm c}_{v}(z)$, then $H_{v}(z^{\prime})\subseteq H_{v}(z)$ and $r^{z^{\prime}}\geq r^{z}$).
$\bullet$  The following definition is of importance.
Definition 1. For a vertex $v$ and a rational assignment $z\in{\cal A}_{v}$, an edge $e\in E_{v}$ is called interesting (in the sense, that an increase at $e$ could promise the existence of an assignment better than $z$) if there exists $z^{\prime}\in{\cal B}_{v}$ such that $z^{\prime}(e)>z(e)$, $z^{\prime}(e^{\prime})=z(e^{\prime})$ for all $e^{\prime}\neq e$, and $C_{v}(z^{\prime})(e)>z(e)$. (In terminology of [1], such an $e$ is called non-satiated.)
For a mixed CF $C_{v}$, one can see that:
  1. (2.5)
    for $z\in{\cal A}_{v}$, an edge $e\in E_{v}$ with $z(e)<b(e)$ is interesting for $v$ under $z$ if and only if $e$ belongs to the tail $T_{v}(z)$.
Indeed, let $z^{\prime}$ be obtained from $z$ by increasing the value at $e$ by a sufficiently small amount $\varepsilon>0$. Then $z^{\prime}\in{\cal B}_{v}$. In case $|z|<q(v)$, we have $e\in T_{v}(z)$, $z^{\prime}\in{\cal A}_{v}$, and $e$ is interesting under $z$ (since $C(z^{\prime})(e)=z^{\prime}(e)>z(e)$). Now let $|z|=q(v)$. If $e\in T_{v}(z)$, then $C_{v}$ does not change $z^{\prime}$ within the tail $T_{v}(z)$ and decreases the values within $H_{v}(z)$ by $\varepsilon/|H_{v}(z)|$ (cf. (2.2)); therefore, $C_{v}(z^{\prime})(e)=z^{\prime}(e)>z(e)$, and $e$ is interesting under $z$ again. And if $e\notin T_{v}(z)$ (i.e., $e$ is in $H_{v}(z)$ or in a tie in $\Pi^{\rm end}_{v}(z)$), then $C_{v}$ applied to $z^{\prime}$ decreases $z^{\prime}(e)$ by $\varepsilon$. So $C(z^{\prime})=z$ and $e$ is not interesting under $z$.
$\bullet$  Now we leave “individual agents” and come to assignments on the whole edge set $E$. An admissible assignment $x\in{\cal B}$ is regarded as rational if it is so for all agents, i.e., $x_{v}\in{\cal A}_{v}$ for all $v\in V$. The set of these is denoted by ${\cal A}$. So any rational assignment $x\in{\cal A}$ respects both the upper bounds $b$ and the quotas $q$.
In what follows an edge connecting vertices $f\in F$ and $w\in W$ is denoted as $fw$.
[…]
We denote the set of stable assignments by ${\cal S}={\cal S}(G,b,q,\geq)$. Using ((2.5)), we can characterize blocking edges for rational assignments as follows:
  1. (2.6)
    for $x\in{\cal A}$, an edge $e=fw\in E$ with $x(e)<b(e)$ is blocking if and only if it belongs to both tails $T_{f}(x)$ and $T_{w}(x)$ (where $T_{v}(x)$ means $T_{v}(x_{v})$).
For stable $x,y\in{\cal S}$, when the relation $x_{f}\succeq y_{f}$ takes place for all firms $f\in F$, we write $x\succeq_{F}y$; similarly, when $x_{w}\succeq y_{w}$ holds for all workers $v\in W$, we write $x\succeq_{W}y$. According to general results on stability for bipartite graphs (subject to the size monotonicity axiom) shown by Alkan and Gale [1], the following is valid:
  1. (2.7)
[…]
    • (a)
      (polarity) the order $\succeq_{F}$ is converse to $\succeq_{W}$, namely: if $x\succeq_{F}y$ then $y\succeq_{W}x$, and vice versa;
    • (b)
      (size invariance) for each fixed $v\in V$, the value $|x_{v}|$ is the same for all $x\in{\cal S}$.
We denote the minimal and maximal elements in $({\cal S},\succeq_{W})$ by $x^{\rm min}$ and $x^{\rm max}$, respectively (so the former is the best, and the latter is the worst w.r.t. $F$).
Let us say that a vertex $v$ is fully filled if the function (restriction) $x_{v}$ is fully filled for $x\in{\cal S}$ (i.e., $|x_{v}|=q(v)$), and deficit otherwise (i.e., $|x_{v}|<q(v)$); this does not depend on the choice of stable $x$, by ((2.7))(b). We will use one more property, strengthening (b) for deficit vertices, which is shown in [1, Corollary 3] for the quota filling case (and therefore, applicable to our mixed model); it says that:
  1. (2.8)
    if $v\in V$ is deficit, then the restriction functions $x_{v}$ are the same for all $x\in{\cal S}$.
3 Finding an initial stable assignment
In order to construct the stable matching $x^{\rm min}$ (which is the best for the part $F$ and the worst for the part $W$) in our mixed problem SMP, one can use two approaches. One of them follows a method applied in [1, Sec. 3.1] to show the existence of a stable assignment in a general case considered there. The other one generalizes ideas of the method of finding the initial stable allocation elaborated in [4].
We prefer to describe the first method, which is shorter and looks more enlightening. It iteratively constructs triples of functions $(b^{i},x^{i},y^{i})$, $i=0,1,\ldots,i,\ldots$  .
In the beginning of the process, one puts $b^{0}:=b$. A current, $i$-th, iteration considers the function $b^{i}$ (which is already known) and proceeds as follows.
At the first stage of the iteration, $b^{i}$ is transformed into $x^{i}$, by putting $x^{i}_{f}:=C_{f}(b^{i}_{f})$ for each $f\in F$. (Then $x^{i}$ satisfies the quotas $q(f)$ for all “firms” $f\in F$, however, it can exceed quotas $q(w)$ for some “workers” $w\in W$.)
At the second stage, the obtained $x^{i}$ is transformed into $y^{i}$, by putting $y^{i}_{w}:=C_{w}(x^{i}_{w})$ for each $w\in W$. (This $y^{i}$ already satisfies the quotas $q(v)$ for all $v\in V$, but it need not be stable.)
The obtained $x^{i}$ and $y^{i}$ are then used to generate the upper bounds $b^{i+1}$ for the next iteration, by putting
$b^{i+1}(e):=\begin{cases}b^{i}(e)&\mbox{if $y^{i}(e)=x^{i}(e)$,}\\ y^{i}(e)&\mbox{if $y^{i}(e)<x^{i}(e)$.}\end{cases}$ (3.1)
Then $b^{i+1}$ is transformed into $x^{i+1}$ and, further, into $y^{i+1}$, and so on.
The process terminates if at some iteration, $i$-th say, we obtain the equality $y^{i}=x^{i}$ (which implies $b^{i+1}=b^{i}$, by (3.1)); then $x^{i}$ is just the required assignment (see the proposition below). In reality, it may happen that the sequence of iterations is infinite. Nevertheless, one can see (by arguing as in [1]) that the process converges. Indeed, the construction shows that for each $i$,
$b^{i}\geq x^{i}\geq y^{i}\qquad\mbox{and}\qquad b^{i}\geq b^{i+1}\geq 0.$ (3.2)
Therefore, $b^{0},b^{1},\ldots$ converge to some function $\widehat{b}\geq 0$. This implies that $x^{0},x^{1},\ldots$ converge to the function $\widehat{x}$, defined by $\widehat{x}_{f}:=C_{f}(\widehat{b}_{f})$ for $f\in F$, and, further, that $y^{0},y^{1},\ldots$ converge to the function $\widehat{y}$ which is equal to $\widehat{x}$ (assuming that all choice functions $C_{v}$ are continuous).
Proposition 3.1
The limit function $\widehat{x}$ is stable and, moreover, $\widehat{x}$ is optimal for $F$, i.e., $\widehat{x}=x^{\rm min}$.
This result was proved by Alkan and Gale for a general case considered in [1]; see Theorem 1 (for the stability) and Theorem 2 (for the optimality) in that paper.
Fortunately, for our mixed model the above method can be modified, by aggregating some sets of consecutive iterations, so as to obtain an efficient algorithm to find the assignment $x^{\rm min}$; its running time is polynomial in $|E|$. We describe this modified method for SMP in the Appendix.
4 Rotations
Throughout this section we consider a fixed stable assignment $x\in{\cal S}$ for our mixed problem SMP. Our aim is to characterize the set of stable assignments $x^{\prime}$ that are close to $x$ and satisfy $x\succ_{F}x^{\prime}$. This will be done by use of a technique of rotations, which generalizes that developed for the stable allocation problem in [4]. First of all we specify some notation needed in what follows.
$\bullet$ We denote by $F^{=}$ and $W^{=}$ the sets of fully filled vertices in the parts $F$ and $W$, respectively, that is, the vertices $f\in F$ and $w\in W$ such that $|x_{f}|=q(v)$ and $|x_{w}|=q(w)$. (By ((2.7))(b), these sets do not depend on the choice of $x\in{\cal S}$. Furthermore, in view of ((2.8)), any transformation $x\rightsquigarrow x^{\prime}\in{\cal S}$ should involve only edges connecting $F^{=}$ and $W^{=}$.)
[…]
$\bullet$ We say that an edge $e\in E$ is saturated by $x$ if $x(e)=b(e)$; the set of such edges is denoted as $U(x)$, and its restriction to $E_{v}$ ($v\in V$) as $U_{v}(x)$.
$\bullet$ Usually we will abbreviate $\succeq_{F}$ to $\succeq$.
We start our description with the observation that property ((2.6)) implies that the stability of $x$ is equivalent to the following property:
  1. (4.1)
    for each non-saturated edge $e=fw\in E-U(x)$, at least one vertex $v$ among $f$ and $w$ is such that: $v$ is fully filled and $e$ belongs to $H_{v}(x)\cup E^{\rm end}_{v}(x)$.
The construction of rotations is based on an analysis of stable assignments close to $x$ as exposed in the next lemma. It uses additional definitions.
Definition 3. Let $f\in F^{=}$. Define the reduced head for $f$ to be the set $\widetilde{H}_{f}=\widetilde{H}_{f}(x)$ consisting of non-saturated edges $e=fw$ in $H_{f}(x)$ such that $e$ belongs to $T_{w}(x)$. For each tie $\pi^{i}$ in $\Pi^{\rm end}_{f}(x)$, define the potential head for $f$ and $\pi^{i}$ to be the set $\widehat{H}_{f}^{i}=\widehat{H}_{f}^{i}(x)$ consisting of non-saturated edges $e=fw$ in $\pi^{i}$ such that $e$ belongs to $T_{w}(x)$, and no edge $e^{\prime}\in E_{f}$ with $e^{\prime}\geq_{f}e$ is incident to a deficit vertex in $W$. The best nonempty set among $\widetilde{H}_{f}(x)$ and $\widehat{H}_{f}^{i}$ ($\pi^{i}\in\Pi^{\rm end}_{f}(x)$) is denoted by $D_{f}=D_{f}(x)$ (ranging the sets according to the order in $\Pi_{f}$); if all these sets are empty, we put $D_{f}:=\emptyset$.
Lemma 4.1
Let $x^{\prime}\in{\cal S}$ be such that $x\succ x^{\prime}$ and $x^{\prime}$ is close to $x$, in the sense that the deviation ($\ell_{1}$-distance) $|x^{\prime}-x|=\sum(|x^{\prime}(e)-x(e)|\colon e\in E)$ is sufficiently small.
  • (i)
    Let $w\in W$ and $x^{\prime}_{w}\neq x_{w}$. Then $w$ is fully filled, and the following are valid:
    • (a)
      $H_{w}(x^{\prime})=H_{w}(x)$, and this is exactly the set of edges $e\in E_{w}$ such that $x^{\prime}(e)<x(e)$;
    • (b)
      no edge in $H_{w}(x)$ is incident to a deficit vertex in $F$.
  • (ii)
    Let $f\in F$ and $x^{\prime}_{f}\neq x_{f}$. Then $f$ is fully filled, and the following are valid:
    • (c)
      if $D_{f}(x)\neq\emptyset$, then $H_{f}(x^{\prime})=D_{f}(x)$, and this is exactly the set of edges $e\in E_{f}$ with $x^{\prime}(e)>x(e)$;
    • (d)
      no edge in $D_{f}$ is incident to a deficit vertex in $W$.
Proof Let $w\in W$ and $x^{\prime}_{w}\neq x_{w}$. Then $w\in W^{=}$ (by ((2.8))). The relation $x\succ_{F}x^{\prime}$ implies $x^{\prime}_{w}\succ_{w}x_{w}$ (by the polarity); therefore, the critical tie for $x^{\prime}_{w}$ is not worse than the one for $x_{w}$: $\pi^{\rm c}_{w}(x^{\prime})\geq\pi^{\rm c}_{w}(x)$ (by ((2.4))). Moreover, since $|x^{\prime}_{w}-x_{w}|=:\varepsilon>0$ is as small as wished, we can conclude that $\pi^{\rm c}_{w}(x^{\prime})=\pi^{\rm c}_{w}(x)$ and $H_{w}(x^{\prime})=H_{w}(x)$, yielding (a). (Here the inclusion $H_{w}(x^{\prime})\supseteq H_{w}(x)$ (valid by ((2.4))) turns into equality since we can choose $\varepsilon>0$ so that any $e\in\pi^{\rm c}_{w}(x)-H_{w}(x)$ and $a\in H_{w}(x)$ satisfy $x^{\prime}(e)-x(e)\leq\varepsilon$, $x(a)-x^{\prime}(a)\leq\varepsilon$, and $x(a)-x(e)>2\varepsilon$, implying $x^{\prime}(a)>x^{\prime}(e)$.) The assertion (b) follows from ((2.8)), taking into account that $x^{\prime}(e)\neq x(e)$ for all $e\in H_{w}(x)$.
Now consider $f\in F$ with $x^{\prime}_{f}\neq x_{f}$. Then $f\in F^{=}$ (by  ((2.8))) and $x_{f}\succ_{f}x^{\prime}_{f}$. Therefore, $\pi^{\rm c}_{f}(x)\geq\pi^{\rm c}_{f}(x^{\prime})$ (by ((2.4))). Suppose that $\pi^{\rm c}_{f}(x)=\pi^{\rm c}_{f}(x^{\prime})$. Then $H_{f}(x^{\prime})\subseteq H_{f}(x)$ (by ((2.4))). For any $e=fw\in H_{f}(x^{\prime})$, we have $x(e)<x^{\prime}(e)\leq b(e)$ (whence $e$ is not saturated by $x$), and $e$ belongs to $T_{w}(x^{\prime})=T_{w}(x)$ (taking into account that $H_{w}(x^{\prime})=H_{w}(x)$ and that $e$ cannot be in $H_{w}(x^{\prime})$). In case $H_{f}(x^{\prime})\neq H_{f}(x)$, any edge $e=fw$ in $H_{f}(x)-H_{f}(x^{\prime})$ belongs $T_{f}(x^{\prime})$, and therefore, $e$ must belong to $H_{w}(x^{\prime})=H_{w}(x)$ (since $x^{\prime}$ is stable). These reasonings imply that $H_{f}(x^{\prime})$ coincides with the reduced head $\widetilde{H}_{f}(x)$, whence $H_{f}(x^{\prime})=D_{f}(x)$ (for terminology and notation, see Definition 3).
Next suppose that $\pi^{\rm c}_{f}(x^{\prime})<\pi^{\rm c}_{f}(x)$. Let $\pi^{\rm c}_{f}(x^{\prime})$ be the tie $\pi^{i}$ in $\Pi_{f}$. Consider an arbitrary edge $e=fw$ in $E^{\rm beg}_{f}(x^{\prime})\cup(\pi^{i}-H_{f}(x^{\prime}))$. This edge belongs to the tail $T_{f}(x^{\prime})$, and since $x^{\prime}$ is stable, $e$ must belong to the head $H_{w}(x^{\prime})=H_{w}(x)$; in particular, the vertex $w$ is not deficit. It follows that $H_{f}(x^{\prime})$ is just the potential head $\widehat{H}^{i}_{f}(x)$ and that all sets $\widetilde{H}_{f}(x)$ and $\widehat{H}^{j}_{f}(x)$, where $\pi^{\rm c}_{f}(x)>\pi^{j}>\pi^{i}$, are empty. Therefore, $H_{f}(x^{\prime})=\widehat{H}^{i}_{f}(x)=D_{f}(x)$, yielding (c). In its turn, (d) follows from the fact that each edge $e=fw\in H_{f}(x^{\prime})$ satisfies $x(e)=0<x^{\prime}(e)$, whence $w$ is not deficit by ((2.8)). $\Box$
When this is not confusing, we will abbreviate notation $H_{v}(x),T_{v}(x),U_{v}(x)$ (where $x$ is fixed) to $H_{v},T_{v},U_{v}$. For $f\in F^{=}$, we call $D_{f}$ defined in (ii) of Lemma 4.1 the active head for $f$ and $x$.
Properties (b) and (d) in Lemma 4.1 can be generalized as follows:
  1. (4.2)
    let $(v_{0},e_{1},v_{1},\ldots,e_{k},v_{k})$ be a path such that: the vertices $v_{0},\ldots,v_{k-1}$ are fully filled while $v_{k}$ is deficit, and for $i=1,\ldots,k$, the edge $e_{k}$ belongs to $H_{v_{i-1}}$ (when $v_{i-1}\in W^{=}$) or $D_{v_{i-1}}$ (when $v_{i-1}\in F^{=}$); then $x^{\prime}(e_{i})=x(e_{i})$ holds for all $i$.
This is shown by induction on $k-i$. Since $v_{k}$ is deficit, we have $x^{\prime}(e_{k})=x(e_{k})$ and, moreover, $x^{\prime}(e)=x(e)$ for all edges $e\in E_{v_{k-1}}$. And for $i=0,\ldots,k-2$, the equality $x^{\prime}(e_{i})=x(e_{i})$ follows from $x^{\prime}(e_{i+1})=x(e_{i+1})$ and the balance condition $\sum(x^{\prime}(a)-x(a)\colon a\in E_{v_{i}})=0$, taking into account that for $a,b\in E_{v_{i}}$, the values $x^{\prime}(a)-x(a)$ and $x^{\prime}(b)-x(b)$ are equal if both $a,b$ belong to the corresponding head for $v_{i}$ and $x^{\prime}$, and have the same sign if both $a,b$ are not in this head.
This property motivates the following Cleaning procedure which is intended for extracting the vertices $v$ in $F^{=}\cup W^{=}$ where no changes of values on $E_{v}$ are possible.
  • (C):
    We extract, step by step, singular vertices in $F^{=}\cup W^{=}$ as follows. In the beginning, we label as singular each vertex $v\in F^{=}\cup W^{=}$ whose head ($D_{v}$ or $H_{v}$) contains an edge incident to a deficit vertex. Also we label as singular each vertex $f\in F^{=}$ such that $D_{f}=\emptyset$ (since this should imply $x^{\prime}_{v}=x_{v}$). At a current step, if there appears a (new) vertex $v\in F^{=}\cup W^{=}$ whose head ($D_{v}$ or $H_{v}$) contains an edge incident to a vertex labeled as singular, then we label $v$ as singular as well.
The vertices in $F^{=}$ and $W^{=}$ not labeled as singular upon termination of this procedure are called regular, and the sets of these are denoted as $F^{+}$ and $W^{+}$, respectively. Note that
  1. (4.3)
    for each $f\in F^{+}$, we have $D_{f}\neq\emptyset$, and all edges $fw\in D_{f}$ are such that $w\in W^{+}$; similarly, for each $w\in W^{+}$, all edges $fw\in H_{w}$ are such that $f\in F^{+}$.
Also by Lemma 4.1(a),(c), the closeness of $x^{\prime}$ to $x$ (subject to $x\succ x^{\prime}$) corresponds to the property of homogeneity, namely:
  1. (4.4)
    for each $f\in F^{+}$, the values $x^{\prime}(e)-x(e)$ are equal to the same number $\phi_{f}\geq 0$ for all edges $e\in D_{f}$; and for each $w\in W^{+}$, the values $x(e)-x^{\prime}(e)$ are equal to the same number $\psi_{w}\geq 0$ for all $e\in H_{w}$.
For $f\in F^{+}$, define $h_{f}:=|D_{f}|$, and for $w\in W^{+}$, define $h_{w}:=|H_{w}|$.
The reasonings above enable us to represent $\phi$ and $\psi$ defined in ((4.4)) via variables in the following system of equalities:
$h_{f}\phi_{f}-\sum(\psi_{w}\colon w\in W^{+},\;fw\in E_{f}-D_{f})=0,\quad f\in F^{+};$ (4.5)
$h_{w}\psi_{w}-\sum(\phi_{f}\colon f\in F^{+},\;fw\in E_{w}-H_{w})=0,\quad w\in W^{+};$ (4.6)
$\phi\geq 0,\;\psi\geq 0.$ (4.7)
(Here $\phi_{f}$ means a uniform increase of $x(e)$ at the edges $e\in D_{f}$, and $\psi_{w}$ a uniform decrease of $x(e)$ at the edges $e\in H_{w}$. The equalities in (4.5) describe the balance conditions for the vertices $f\in F^{+}$, and the ones in (4.6) describe similar conditions for the vertices $w\in W^{+}$, keeping the corresponding quota equalities.)
Also we should add to the above relations the following equality:
$\sum(h_{f}\phi_{f}\colon f\in F^{+})=\varepsilon$ (4.8)
for a sufficiently small number $\varepsilon>0$ (in order to avoid the equality $x^{\prime}=x$).
We fix such an $\varepsilon$, and our aim is to show that the set (polytope)
${\cal P}={\cal P}^{\varepsilon}(x):=\{(\phi,\psi)\}\subset{\mathbb{R}}^{F^{+}}\times{\mathbb{R}}^{W^{+}},$ (4.9)
where $(\phi,\psi)$ range the solutions to system (4.5)–(4.8), matches the set of augmentations $x^{\prime}-x$ for all $x^{\prime}\in{\cal S}$ such that $x^{\prime}\prec x$ and $x^{\prime}$ is close to $x$ (measured by the parameter $\varepsilon$).
More precisely, let us associate with $(\phi,\psi)\in{\cal P}$ the following vector $\Delta^{\phi,\psi}\in{\mathbb{R}}^{E}$:
  1. (4.10)
    for $e=fw\in E$, define $\Delta^{\phi,\psi}(e)$ to be equal to $\phi_{f}$ when $f\in F^{+}$ and $e\in D_{f}$, to be $-\psi_{j}$ when $w\in W^{+}$ and $e\in H_{w}$, and 0 otherwise.
Also define the set
$X^{\varepsilon}(x):=\{x+\Delta^{\phi,\psi}\colon(\phi,\psi)\in{\cal P}\}.$ (4.11)
Theorem 4.2
Let $x\in{\cal S}-\{x^{\rm max}\}$. Then for a sufficiently small $\varepsilon>0$,
(a) $X^{\varepsilon}(x)$ is exactly the set of stable assignments $x^{\prime}\in{\cal S}$ such that $x^{\prime}\prec x$ and the deviation $\sum(|x^{\prime}(e)-x(e)|\colon e\in E)$ is equal to $2\varepsilon$;
(b) if $y\in{\cal S}$ and $y\prec x$, then there exists $x^{\prime}\in X^{\varepsilon}(x)$ satisfying $y\preceq x^{\prime}\prec x$.
(Also we will explain that system (4.5)–(4.8) has no solution if and only if $x=x^{\rm max}$.)
The proof of this theorem is given in the rest of this section. One auxiliary assertion that we use in the proof, which is interesting in its own right, is as follows.
Lemma 4.3
Let $A$ and $B$ be two disjoint nonempty subsets of $E$, and let $(J,A\cup B)$ be the subgraph of $G$ induced by $A\cup B$. For a vertex $v\in J$, let $A_{v}$ ($B_{v}$) denote the set of edges in $A$ (resp. $B$) incident to $v$. Suppose that
  1. (4.12)
    for each edge $fw\in A$, the set $B_{w}$ is nonempty, and similarly, for each $fw\in B$, the set $A_{f}$ is nonempty.
Then there exist nonzero functions $\alpha:A\to{\mathbb{R}}_{+}$ and $\beta:B\to{\mathbb{R}}_{+}$ such that:
(i) for each $v\in J$, the value $\alpha(A_{v})-\beta(B_{v})$ is zero;
(ii) for each $f\in J\cap F$, the values $\alpha(e)$ for all edges $e$ in $A_{f}$ are equal, and similarly, for each $w\in J\cap W$, the values $\beta(e)$ for all $e\in B_{w}$ are equal.
Proof Let $J_{F}:=J\cap F$ and $J_{W}:=J\cap W$. For $a\in{\mathbb{R}}_{+}^{J_{F}}$, denote by $\alpha^{a}$ the function on $A$ taking the value $a(f)$ for each edge $fw\in A_{f}$. For $b\in{\mathbb{R}}_{+}^{J_{W}}$, denote by $\beta^{b}$ the function on $B$ taking the value $b(w)$ for each edge $fw\in B_{w}$.
Define operators $p:{\mathbb{R}}_{+}^{J_{F}}\to{\mathbb{R}}_{+}^{J_{W}}$ and $q:{\mathbb{R}}^{J_{W}}\to{\mathbb{R}}_{+}^{J_{F}}$ as follows:
  1. (4.13)
    • (a)
      for $a\in{\mathbb{R}}_{+}^{J_{F}}$ and $w\in J_{W}$, put $p(a)_{w}:=\sum(a(f)\colon fw\in A_{w})\,/\,|B_{w}|$;
    • (b)
      for $b\in{\mathbb{R}}_{+}^{J_{W}}$ and $f\in J_{F}$, put $q(b)_{f}:=\sum(b(w)\colon fw\in B_{f})\,/\,|A_{f}|$;
and let $r$ be their composition $q\circ p$ (which maps ${\mathbb{R}}_{+}^{J_{F}}$ into itself). These $p,q,r$ are well defined since condition ((4.12)) implies that for each $f\in J_{F}$, $A_{f}\neq\emptyset$, and similarly for each $w\in J_{W}$, $B_{w}\neq\emptyset$ (though some $A_{w}$ ($w\in J_{W}$) or $B_{f}$ ($f\in J_{f}$) may be empty).
Fix a number $\xi>0$, and consider the set $Q$ formed by the vectors $a\in{\mathbb{R}}_{+}^{J_{F}}$ such that $\sum(a(f)|A_{f}|\colon f\in J_{F})=\xi$. This set forms a nonempty polytope (taking into account that each $a(f)$ is nonnegative and bounded, in view of $A_{f}\neq\emptyset$).
Now using ((4.13))(a), for $a\in{\mathbb{R}}_{+}^{J_{F}}$ and $b:=p(a)$, we obtain
$\sum(b(w)|B_{w}|\colon w\in J_{W})=\sum\nolimits_{w\in J_{W}}\sum(a(f)\colon fw\in A_{w})=\sum(a(f)|A_{f}|\colon f\in J_{F}).\quad$
Similarly, ((4.13))(b) implies that for $b\in{\mathbb{R}}_{+}^{J_{W}}$ and $a^{\prime}:=q(b)$,
$\sum(a^{\prime}(f)|A_{f}|\colon f\in J_{F})=\sum\nolimits_{f\in J_{F}}\sum(b(w)\colon fw\in B_{f})=\sum(b(w)|B_{w}|\colon w\in J_{W}).$
It follows that for $a\in Q$, the vector $a^{\prime}:=r(a)=q(p(a))$ belongs to $Q$ as well. So the operator $r$ maps $Q$ into itself. Since $Q$ is a polytope and $r$ is, obviously, continuous, we can apply Brouwer fixed point theorem and conclude that there exists $a\in Q$ such that $r(a)=a$. Then the functions $\alpha=\alpha^{a}$ and $\beta=\beta^{p(a)}$ are as required in (i)–(ii). $\Box$
Definition 4. Let us call an edge $e=fw$ active for $F$ if $f\in F^{+}$ and $e\in D_{f}$, and active for $W$ if $w\in W^{+}$ and $e\in H_{w}$. The subgraph of $G$ induced by the set of active edges is called the active graph and denoted as $\Gamma=(V_{\Gamma},E_{\Gamma})$. The directed counterpart of the active graph, denoted as $\overrightarrow{\Gamma}=(V_{\Gamma},\overrightarrow{E}_{\Gamma})$, is obtained by replacing each edge $e=fw\in E_{\Gamma}$ by the directed edge $\overrightarrow{e}=(f,w)$ if $e\in D_{f}$, and by the directed edge $\overrightarrow{e}=(w,f)$ if $e\in H_{w}$. A (weak) component of $\Gamma$ (or $\overrightarrow{\Gamma}$) is a maximal subgraph $K$ in it such that any two vertices of $K$ are connected by a (not necessarily directed) path.
Using Lemma 4.3, one can obtain the following
Lemma 4.4
Let $K=(V_{K},E_{K})$ be a component of the active graph $\Gamma$. Then (4.5)–(4.8) has a solution $(\phi,\psi)$ taking zero values on all vertices outside $K$.
Proof  Take as $A$ the set $\cup(D_{f}\colon f\in F_{K})$ and take as $B$ the set $\cup(H_{w}\colon w\in W_{K})$, where $F_{K}:=F\cap V_{K}$ and $W_{K}:=W\cap V_{K}$. Property ((4.3)) implies that $A,B$ obey condition ((4.12)). Then, by Lemma 4.3, there exist $\alpha:A\to{\mathbb{R}}_{+}$ and $\beta:B\to{\mathbb{R}}_{+}$ satisfying properties (i),(ii) in that lemma; moreover, as is seen from its proof, for any number $\xi>0$, the functions $\alpha,\beta$ can be chosen so as to satisfy the equality $\sum(\alpha(e)\colon e\in A)=\xi$. Take $\xi:=\varepsilon$, define $\phi_{f}:=\alpha(e)$ for $f\in F_{K}$ and $e\in D_{f}$, and define $\psi_{w}:=\beta(e)$ for $w\in W_{K}$ and $e\in H_{w}$. These $\phi,\psi$ extended by zero outside $K$ satisfy (4.5)–(4.8), as required. $\Box$
Lemma 4.4 has a sharper version. To state it, consider the directed active graph $\overrightarrow{\Gamma}=(V_{\Gamma},\overrightarrow{E}_{\Gamma})$ (see Definition 4) and its strong components. (Recall that a strong component of a directed graph is a maximal subgraph $L=(V_{L},E_{L})$ of this graph such that $|V_{L}|>1$ and for any ordered pair $(u,v)$ in $V_{L}$, there is a directed path from $u$ to $v$. In other words, $L$ is a maximal connected subgraph representable as the union of directed cycles.) A weak component $K$ in $\overrightarrow{\Gamma}$ may contain several strong components; we denote the set of these by ${\cal L}(K)$. There is a natural partial order on ${\cal L}(K)$ in which $L\in{\cal L}(K)$ is regarded as preceding $L^{\prime}\in{\cal L}(K)$ if $K$ has a directed path going from $L$ to $L^{\prime}$. We denote by ${\cal L}^{\rm max}(K)$ the set of maximal components in ${\cal L}(K)$ (i.e., those having no succeeding components). Note that ((4.3)) implies that if $v$ is a vertex of $L\in{\cal L}^{\rm max}(K)$, then any edge of $K$ leaving $v$ belongs to $L$ as well.
We associate with $(\phi,\psi)\in{\mathbb{R}}_{+}^{F^{+}}\times{\mathbb{R}}_{+}^{W^{+}}$ the function $\overrightarrow{\Delta}=\overrightarrow{\Delta}^{\phi,\psi}$ on the set of active directed edges $\overrightarrow{E}_{\Gamma}$, taking the value $\phi_{f}$ on the edges $(f,w)\in D_{f}$, and $\psi_{w}$ on the edges $(w,f)\in H_{w}$ (cf. $\Delta^{\phi,\psi}$ defined in ((4.10))). This $\overrightarrow{\Delta}$ represents a circulation in $\overrightarrow{\Gamma}$, because the equalities in (4.5) are equivalent to the conservation (or balance) conditions $\overrightarrow{\Delta}(\delta^{\rm out}(f))=\overrightarrow{\Delta}(\delta^{\rm in}(f))$ for $f\in F^{+}$, and similarly for (4.6) and $W^{+}$. Here $\delta^{\rm in}(v)$ ($\delta^{\rm out}(v)$) denotes the set of edges in $\overrightarrow{\Gamma}$ entering (resp. leaving) a vertex $v$. The following fact is of use:
  1. (4.14)
    for $(\phi,\psi)\in{\cal P}$, if vertices $u,v$ are connected in $\overrightarrow{\Gamma}$ by a directed path from $u$ to $v$, and if $(\phi,\psi)_{u}>0$, then $(\phi,\psi)_{v}>0$ as well.
(It suffices to see this for a single edge $e=(u,v)$ in $\overrightarrow{\Gamma}$. This edge belongs to the head at $u$ (either $H_{u}$ or $D_{u}$, in the directed version), and if $(\phi,\psi)_{u}>0$, then $\overrightarrow{\Delta}^{\phi,\psi}(e)>0$, implying $(\phi,\psi)_{v}>0$, by the balance condition at $v$.)
Now Lemma 4.4 is strengthened as follows.
Lemma 4.5
Let $K$ be a weak component of $\overrightarrow{\Gamma}$. Then for any maximal component $L\in{\cal L}^{\rm max}(K)$, there exists a solution $(\phi,\psi)\in{\mathbb{R}}_{+}^{F^{+}}\times{\mathbb{R}}_{+}^{W^{+}}$ to system (4.5)–(4.8) whose values outside $L$ are zeros, and this solution is unique.
Proof  Consider $L=(V_{L},E_{L})\in{\cal L}^{\rm max}(K)$. As is mentioned above, for each vertex $v\in V_{L}$, the set $\delta^{\rm out}(v)$ of edges of $\overrightarrow{\Gamma}$ leaving $v$ is contained in $L$. So we can apply Lemma 4.3 to the sets $A:=\cup(D_{f}\colon f\in F_{L})$ and $B:=\cup(H_{w}\colon w\in W_{L})$. This gives rise to a solution $(\phi,\psi)\in{\cal P}$ to (4.5)–(4.8) having the support $\{v\colon(\phi,\psi)_{v}\neq 0\}$ within $L$.
The see the uniqueness of $(\phi,\psi)$, we first observe that its support is exactly $V_{L}$. (This follows for the strong connectedness of $L$ and property ((4.14)).)
Now suppose, for a contradiction, that there is another solution $(\phi^{\prime},\psi^{\prime})\in{\cal P}$ having the support within $L$. Then one can choose numbers $\lambda,\mu>0$ such that the function $(\phi^{\prime\prime},\psi^{\prime\prime}):=(\lambda\phi,\lambda\psi)-(\mu\phi^{\prime},\mu\psi^{\prime})$ is nonnegative, satisfies (4.5)–(4.8), and takes zero value on some $f\in F_{L}$ or $w\in W_{L}$. But this is impossible by the observation above. $\Box$
On the other hand, one can show the following
Lemma 4.6
For each weak component $K$ of $\overrightarrow{\Gamma}$, any function $(\phi,\psi)\in{\mathbb{R}}_{+}^{F^{+}}\times{\mathbb{R}}_{+}^{W^{+}}$ satisfying (4.5)–(4.7) takes zero value on all vertices that belong to none of the maximal strong components in $K$.
Proof Let $Y$ be the set of vertices occurring in components in ${\cal L}^{\rm max}(K)$. Consider a function $(\phi,\psi)$ satisfying (4.5)–(4.7), and suppose, for a contradiction, that the set $Z$ of vertices $v\in V_{K}-Y$ with nonzero values $(\phi,\psi)_{v}$ is nonempty. Since the strong components in ${\cal L}^{\rm max}(K)$ are maximal, the set $\delta^{\rm out}(Y)$ of edges of $K$ going from $Y$ to $V_{K}-Y$ is empty.
On the other hand, for any vertex $u$ in $V_{K}-Y$, in particular, for $u\in Z$, there is a directed path $P$ going from $u$ to a maximal strong component in $K$. Then, by ((4.14)), $(\phi,\psi)_{u}>0$ implies $(\phi,\psi)_{v}>0$ for all vertices $v$ on this path; therefore, the edge $e$ of $P$ leaving $V_{K}-Y$ and entering $Y$ satisfies $\overrightarrow{\Delta}^{\phi,\psi}(e)>0$.
As a consequence, we obtain
$\overrightarrow{\Delta}(\delta^{\rm in}(Y))-\overrightarrow{\Delta}(\delta^{\rm out}(Y))=\overrightarrow{\Delta}(\delta^{\rm in}(Y))>0,$
where $\delta^{\rm in}(Y)$ is the set of edges of $K$ entering $Y$. But the difference in the left hand side must be zero (as this is the sum, over the vertices $v$ of $Y$, of “divergences” $\overrightarrow{\Delta}(\delta^{\rm in}(v))-\overrightarrow{\Delta}(\delta^{\rm out}(v))$, which are zeros). A contradiction. $\Box$
Let ${\cal L}^{\rm rot}={\cal L}^{\rm rot}(x)$ denote the set of maximal strong components $L$ occurring in the sets ${\cal L}^{\rm max}(K)$ over all weak components $K$ of the active graph $\overrightarrow{\Gamma}(x)$. By Lemma 4.5, for each component $L\in{\cal L}^{\rm rot}$, there is a unique solution to (4.5)–(4.8) having the support in $L$ (which is, moreover, exactly $V_{L}$); we denote this solution as $(\phi_{L}^{\varepsilon},\psi_{L}^{\varepsilon})=(\phi_{L}^{\varepsilon}(x),\psi_{L}^{\varepsilon}(x))$. Then Lemmas 4.5 and 4.6 imply
Corollary 4.7
The vertices of the polytope ${\cal P}^{\varepsilon}(x)$ (defined in (4.9)) are exactly the vectors $(\phi_{L}^{\varepsilon},\psi_{L}^{\varepsilon})$ for $L\in{\cal L}^{\rm rot}$. The supports of these vectors are pairwise disjoint.
This enables us to specify the notion of rotations in our model.
Definition 5. By a ${\mathbb{Z}}$-circulation we mean an integer-valued function $\rho:E\to{\mathbb{Z}}$ such that $\rho\neq 0$ and for each vertex $v\in V$, the sum $\rho(E_{v})$ ($=\sum(\rho(e)\colon e\in E_{v})$) is zero. Such a $\rho$ is called aligned if $\rho(fw),\rho(fw^{\prime})>0$ implies $\rho(fw)=\rho(fw^{\prime})$, and, symmetrically, $\rho(fw),\rho(f^{\prime}w)<0$ implies $\rho(fw)=\rho(f^{\prime}w)$. We say that $\rho$ is normalized if the g.c.d. of the values $\rho(e)$, $e\in E$, is one. We refer to a normalized aligned ${\mathbb{Z}}$-circulation $\rho$ for which the subgraph of $G$ induced by the support $\{e\in E\colon\rho(e)\neq 0\}$ of $\rho$ is connected as a pre-rotation in $G$.
Definition 6. Consider a maximal strong component $L\in{\cal L}^{\rm rot}$ and the corresponding vector $(\phi_{L}^{\varepsilon},\psi_{L}^{\varepsilon})\in{\cal P}^{\varepsilon}(x)$ with the support $V_{L}$ (existing by Lemma 4.5). Then there is a normalized aligned $Z$-circulation (viz. pre-rotation) proportional to the function $\Delta=\Delta^{\phi_{L}^{\varepsilon},\psi_{L}^{\varepsilon}}$ on $E$ defined as in ((4.10)); we denote this circulation by $\rho_{L}$ and call it the rotation for $x$ associated with $L$. Its analog for $\overrightarrow{\Gamma}$ is denoted as $\overrightarrow{\rho}_{L}$ (which for $\overrightarrow{e}\in\overrightarrow{E}_{\Gamma}$, takes the value $|\Delta(e)|$). We say that a weight $\lambda\in{\mathbb{R}}_{+}$ is admissible for $L$, or for $\rho_{L}$, under the assignment $x$ if the following holds:
  1. (4.15)
    (a) the function $x^{\prime}:=x+\lambda\rho_{L}$ is nonnegative and does not exceed the upper bound $b$, and (b) for each $w\in W^{+}$, either $\pi^{\rm c}_{w}(x^{\prime})=\pi^{\rm c}_{w}(x)$ and $H_{w}(x^{\prime})\supseteq H_{w}(x)$, or $x^{\prime}(e)=0$ for all edges $e$ in $\pi^{\rm c}_{w}(x)$ (in the latter case, $x(e)=\lambda$ is valid, and the critical tie at $w$ becomes better: $\pi^{\rm c}_{w}(x^{\prime})>\pi^{\rm c}_{w}(x)$).
The maximal admissible weight $\lambda$ for $L$ is denoted by $\tau_{L}=\tau_{L}(x)$ and called the maximal rotational weight for $L$, or for $\rho_{L}$.
Remark 1. It is well known that all rotations arising in stable allocation problems are $0,\pm 1$ functions whose supports are formed by simple cycles. In contrast, in SMP, rotations may take large values (even exponentially large in $|E|$). An instance of a “big” rotation in a “small” graph will be demonstrated in Example 3 in Sect. 5.
Note that Lemma 4.5 provides the existence of an admissible weight $\lambda>0$ for each $L\in{\cal L}^{\rm rot}$. Clearly if $\lambda$ is admissible, then any $\lambda^{\prime}$ in the interval $[0,\lambda]$ is admissible as well. Moreover, as a refinement of ((4.15)) in case $\lambda<\tau_{L}$, one can see that
  1. (4.16)
    if $\lambda<\tau_{L}$, then the transformation of $x$ into $x^{\prime}:=x+\lambda\rho_{L}$ preserves the active graph $\Gamma$ and: for each $w\in V^{+}$, we have $\pi^{\rm c}_{w}(x^{\prime})=\pi^{\rm c}_{w}(x)$ and $H_{w}(x^{\prime})=H_{w}(x)$, and for each $f\in F^{+}$, we have $H_{f}(x^{\prime})=D_{f}(x)$.
This implies that in case $\lambda<\tau_{L}$, the rotation $\rho_{L}$ remains applicable to the new $x^{\prime}$, but now the maximal weight for $\rho_{L}$ reduces to $\tau_{L}(x)-\lambda$ ($=\tau_{L}(x^{\prime})$).
Next, the importance of admissible weights is emphasized by the following
Lemma 4.8
For any $0<\lambda\leq\tau_{L}$, the function $x^{\prime}:=x+\lambda\Delta_{L}$ is a stable assignment, and $x^{\prime}\prec x$.
Proof  We have $0\leq x^{\prime}\leq b$ (by ((4.15))), and $|x^{\prime}_{v}|=|x_{v}|$ for all $v\in V$ (by the definition of $\rho_{L}$). So $x^{\prime}$ is a rational assignment with the same sets $F^{=}$ and $W^{=}$.
To show the stability of $x^{\prime}$, consider an edge $e=fw\in E$ with $x^{\prime}(e)<b(e)$. We have to show that $e$ is not blocking for $x^{\prime}$, i.e., there is $v\in\{f,w\}$ such that the edge $e$ either belongs to the head $H_{v}(x^{\prime})$ or occurs in a tie in $\Pi^{\rm end}_{v}(x^{\prime})$; then $e$ is not interesting for $v$ under $x^{\prime}$, and we are done. (Here it suffices to consider the case when both $f,w$ are fully filled; for otherwise Cleaning procedure ensures that $x^{\prime}_{f}=x_{f}$ and $x^{\prime}_{w}=x_{w}$.)
To show this, we first suppose that $x^{\prime}(e)\neq x(e)$. The case $x^{\prime}(e)<x(e)$ is possible only if $w\in V_{L}$ and $e\in H_{w}(x)$. Then (cf. ((4.15))(b)) either $e\in H_{w}(x^{\prime})$ or $e\in\pi^{i}<\pi^{\rm c}_{w}(x^{\prime})$. And if $x^{\prime}(e)>x(e)$, then $f\in V_{L}$ and $e\in D_{f}(x)$, implying $e\in H_{f}(x^{\prime})$. Hence in both cases, $e$ is not blocking for $x^{\prime}$.
Now let $x^{\prime}(e)=x(e)<b(e)$. Then $e\notin E_{L}$. Since $e$ is not blocking for $x$ (as $x$ is stable), we are in at least one of the following cases: (i) $e\in H_{w}$; (ii) $e\in E^{\rm end}_{w}(x)$; (iii) $e\in H_{f}(x)$; (iv) $e\in E^{\rm end}_{f}(x)$ (where $w\in W^{=}$ in cases (i),(ii), and $f\in F^{=}$ in cases (iii),(iv)).
Note that if none of $f,w$ is in $V_{L}$, then $x^{\prime}_{f}=x_{f}$ and $x^{\prime}_{w}=x_{w}$, and the result follows. So assume that some of $f,w$ is in $V_{L}$ (while $e\notin E_{L}$). Moreover, we may assume that $w\in V_{L}$ in cases (i),(ii); for otherwise $x^{\prime}_{w}=x_{w}$, and $e$ is not interesting for $w$ under $x^{\prime}$ (since this is so under $x$). Similarly, we may assume that $f\in V_{L}$ in cases (iii),(iv). Now consider these cases in detail.
In case (i) with $w\in V_{L}$, we have $e\in E_{L}$; so this is not the case.
In case (ii) with $w\in V_{L}$, we have $\pi^{\rm c}_{w}(x^{\prime})\geq\pi^{\rm c}_{w}(x)$. Then $e\in\pi^{i}<\pi^{\rm c}_{w}(x)$ implies $e\in\pi^{i}<\pi^{\rm c}_{w}(x^{\prime})$, whence $e$ is not interesting for $w$ under $x^{\prime}$.
In case (iii) with $f\in V_{L}$, we have $e\in H_{f}(x)$ but $e\notin\widetilde{H}_{f}(x)$ (for otherwise there would be $e\in\widetilde{H}_{f}(x)=D_{f}(x)$, implying $x^{\prime}(e)>x(e)$). This implies that $e\in H_{w}(x)$. (Here we appeal to Definition 3 and notice that $w$ cannot be deficit, for otherwise $f$ would be labeled as singular by Cleaning procedure.) Then we are as in case (i).
Finally, in case (iv) with $f\in V_{L}$, compare $\pi^{i}\in\Pi_{f}$ (containing $e$) with $\pi^{\rm c}_{f}(x^{\prime})=\pi^{j}$ (containing $H_{f}(x^{\prime})=D_{f}(x)$). If $\pi^{i}<\pi^{j}$, we are done. So assume that $\pi^{i}\geq\pi^{j}$. The vertex $w$ is not deficit. Also $e\notin D_{f}(x)$ (otherwise we obtain $x^{\prime}(e)>x(e)$). It follows (cf. Definition 3) that the critical tie $\pi^{\rm c}_{w}(x)$ must be better than the tie in $\Pi_{w}$ containing $e$. This implies that $e\in E^{\rm end}_{w}(x^{\prime})$, and $e$ is not interesting for $w$ under $x^{\prime}$.
Thus, in all cases, $e$ is not blocking for $x^{\prime}$, as required. The fact $x^{\prime}\prec x$ is easy. $\Box$
In fact, different rotations in ${\cal R}(x):=\{\rho_{L}\colon L\in{\cal L}^{\rm rot}(x)\}$ can be applied independently, and the following generalization of Lemma 4.8 is valid.
Proposition 4.9
Let $\lambda:{\cal L}^{\rm rot}(x)\to{\mathbb{R}}_{+}$ satisfy $\lambda(L)\leq\tau_{L}$ for all $L\in{\cal L}^{\rm rot}(x)$. Then $x^{\prime}:=x+\sum(\lambda(L)\Delta_{L}\colon L\in{\cal L}^{\rm rot}(x))$ is a stable assignment.
Proof This is shown by use of Lemma 4.8 and the fact that the maximal strong components of the active graph are pairwise disjoint (the latter implies, in particular, that $x^{\prime}$ does not exceed upper bounds and quotas). More precisely, we have to show that each edge $e=fw\in E$ is non-blocking for $x^{\prime}$. This is immediate if both vertices $f,w$ belong to the same component in ${\cal L}^{\rm rot}$ (by appealing to Lemma 4.8), or if both equalities $x^{\prime}_{f}=x_{f}$ and $x^{\prime}_{w}=x_{w}$ take place (since the stability of $x$ implies that $e$ is not blocking for $x$ and $x^{\prime}$).
So we may assume that $e$ belongs to no component in ${\cal L}^{\rm rot}$ (implying $x^{\prime}(e)=x(e)$) and that either $x^{\prime}_{f}\neq x_{f}$ or $x^{\prime}_{w}\neq x_{w}$ (or both) takes place.
Arguing as in the proof of Lemma 4.8, we get into at least one of cases (i)–(iv) examined there; moreover, we have $w\in V_{L}$ in case (i) or (ii), and $f\in V_{L}$ in case (iii) or (iv), where $L\in{\cal L}^{\rm rot}$. In cases (i),(ii) (when $w\in V_{L}$), our analysis is similar to that in the previous proof. And in cases (iii),(iv) (when $f\in V_{L}$), if the other end $w$ of $e$ does not belong to any other component in ${\cal L}^{\rm rot}$, then we again argue as in the previous proof. And if $w$ belongs to another component $L^{\prime}\in{\cal L}^{\rm rot}$, then we can switch to consideration of this $L^{\prime}$ rather than $L$, where we get into case (i) or (ii). $\Box$
As a consequence of this proposition, we obtain the following
Corollary 4.10
For a sufficiently small $\varepsilon>0$, the set $X^{\varepsilon}(x)$ defined in (4.11) consists of stable assignments.
This gives one direction in assertion (a) of Theorem 4.2. The other (easier) direction, saying that any stable assignment $x^{\prime}\in{\cal S}$ close to $x$ and satisfying $x^{\prime}\prec x$ belongs to $X^{\varepsilon}(x)$ for an appropriate $\varepsilon$, can be concluded from our discussion in the beginning of this section (including Lemma 4.1 and Cleaning procedure).
Now we show part (b) of this theorem (which leads to the nice property that any stable $y\prec x$ can be reached from $x$ by a sequence of rotations).
Lemma 4.11
If $x,y\in{\cal S}$ are such that $y\prec x$, then for a sufficiently small $\varepsilon>0$, there exists $x^{\prime}\in X^{\varepsilon}(x)$ between $x$ and $y$, namely, $y\preceq x^{\prime}\prec x$.
Proof Define $E^{>}:=\{e\in E\colon y(e)>x(e)\}$ and $E^{<}:=\{e\in E\colon y(e)<x(e)\}$. For $v\in V$, define $E^{>}_{v}:=E^{>}\cap E_{v}$ and $E^{<}_{v}:=E^{<}\cap E_{v}$. Clearly each edge $fw$ in $E^{>}\cup E^{<}$ connects fully filled vertices: $f\in F^{=}$ and $w\in W^{=}$. Let $V^{\gtrless}$ denote the set of vertices covered by the edges in $E^{>}$ (and $E^{<}$).
First of all we specify the structure and interrelation of $E^{>}_{v}$ and $E^{<}_{v}$ for $v\in V^{\gtrless}$. We observe the following:
  1. (4.17)
    for $f\in V^{\gtrless}\cap F$, the relation $C_{f}(x_{f}\vee y_{f})=x_{f}$ (in view of $x_{f}\succ y_{f}$) implies $E^{>}_{f}\subseteq H_{f}(x)\cup E^{\rm end}_{f}(x)$ and $E^{<}_{f}\subseteq E^{\rm beg}_{f}(x)\cup\{\pi^{\rm c}_{f}(x)\}$; and
  1. (4.18)
    for $w\in V^{\gtrless}\cap W$, the relation $C_{w}(x_{w}\vee y_{w})=y_{w}$ (in view of $y\succ_{W}x$) implies that either $E^{>}_{w}\subseteq T_{w}(x)$ and $H_{w}(x)\subseteq E^{<}_{w}\subseteq\pi^{\rm c}_{w}(x)$ (when $\pi^{\rm c}_{w}(y)=\pi^{\rm c}_{w}(x)$), or $E^{>}_{w}\subseteq E^{\rm beg}_{w}(x)$ and $E^{<}_{w}=\{e\in\pi^{\rm c}_{w}(x)\colon x(e)>0\}$ (when $\pi^{\rm c}_{w}(y)>\pi^{\rm c}_{w}(x)$).
Note that from there relations one can conclude that
  1. (4.19)
    for any $f\in V^{\gtrless}\cap F$, there holds $E^{>}_{f}\subseteq\widetilde{H}_{f}(x)\cup\cup(\widehat{H}^{i}_{f}(x)\colon\pi^{i}\in\Pi^{\rm end}_{f}(x))$ (where $\widetilde{H}_{f}$ and $\widehat{H}^{i}_{f}$ are defined in Definition 3).
Using these observations, we construct sets $A_{v}$ of active edges for $x$ and $v\in V^{\gtrless}$ by use of the following procedure. Initially, we fix an arbitrary vertex $f=f_{0}$ in $V^{\gtrless}\cap F$. Relation ((4.19)) implies that $D_{f}(x)\neq\emptyset$. Define $A_{f}:=D_{f}(x)$ and $W(A_{f}):=\{w\colon fw\in A_{f}\}$. We assert that
$W(A_{f})\subseteq V^{\gtrless}\cap W.$ (4.20)
[…]
Now using (4.20), we proceed as follows. For each $w\in W(A_{f})$, assign $A_{w}:=H_{w}(x)$. Then $A_{w}\subseteq E^{<}_{w}$ (cf. ((4.18))). For each $e=fw\in A_{w}$, the vertex $f$ belongs to $V^{\gtrless}$ (since $e\in E^{<}$), and we handle $f$ in a way similar to that applied to $f_{0}$ above. And so on.
As a result, this procedure constructs a collection ${\cal A}$ of sets $A_{v}$ for certain vertices $v$ in $V^{\gtrless}$. Here if $v=f\in F$, then $A_{f}$ coincides with $D_{f}(x)$, and if $v=w\in W$, then $A_{w}=H_{w}$. Upon termination of the procedure, for each edge $fw\in A_{f}\in{\cal A}$, the collection ${\cal A}$ contains the set $A_{w}$, and similarly, for each edge $fw\in A_{w}\in{\cal A}$, ${\cal A}$ contains $A_{f}$. In other words, ${\cal A}$ determines a subgraph $\overrightarrow{\Gamma}^{\prime}$ of the active graph $\overrightarrow{\Gamma}(x)$, which is closed under the reachability by directed paths. Then $\overrightarrow{\Gamma}^{\prime}$ contains a maximal strong component $L$ of $\overrightarrow{\Gamma}(x)$. We fix $L$, take the rotation $\rho_{L}$ along with a sufficiently small weight $\lambda>0$, and define $x^{\prime}:=x+\lambda\rho_{L}$. We assert that this $x^{\prime}$ is as required.
To see this, we first observe that by the construction and explanations above,
  1. (4.21)
    for each $w\in V_{L}\cap W$, we have $A_{w}=H_{w}\subseteq E^{<}_{w}$, and $y(e)<x^{\prime}(e)<x(e)$ holds for all $e\in A_{w}$.
In its turn, for each $f\in V_{L}\cap F$, the edges $e\in A_{f}$ satisfy $x(e)<x^{\prime}(e)<b(e)$. We need somewhat stronger relations involving $y$. More precisely, we assert that
  1. (4.22)
    for each $f\in V_{L}\cap F$, any edge $e=fw\in A_{f}$ ($=D_{f}(x)$) with $y(e)<b(e)$ belongs to either $E^{>}\cup H_{w}(y)$ or to a tie in $\Pi^{\rm end}_{w}(y)$.
Indeed, since $e$ is not blocking for $y$, $e$ belongs to at most one tail among $T_{f}(y)$ and $T_{w}(y)$ (cf.((2.6))); therefore, at least one of the following takes place: (i) $e\in H_{f}(y)$; (ii) $e\in H_{w}(y)$; (iii) $e$ belongs to a tie in $\Pi^{\rm end}_{f}(y)$; and (iv) $e$ belongs to a tie in $\Pi^{\rm end}_{w}(y)$.
In cases (ii) and (iv), we are done. Assume that we are not in these cases. Then $e\in T_{w}(y)$. Also, by the definition of $D_{f}(x)$, we have $e\in T_{w}(x)$. Let $\pi^{i}$ and $\pi^{j}$ be the ties in $\Pi_{f}$ containing $D_{f}(x)$ and $H_{f}(y)$, respectively. In case (iii), $\pi^{j}$ is better than $\pi^{i}$ (containing $e$), which is impossible. Finally, in case (i), we have $e\in D_{f}(x)\cap H_{f}(y)$, whence $i=j$. Suppose that $\pi^{i}=\pi^{\rm c}_{f}(x)$. Then $y_{f}\preceq x_{f}$ and $y_{f}\neq x_{f}$ (in view of $f\in V^{\gtrless}$) imply that $H_{f}(y)\subseteq H_{f}(x)$ and $H_{f}(y)\subseteq E^{>}$, whence $e\in E^{>}$. And if $\pi^{i}$ is worse than $\pi^{\rm c}_{f}(x)$, then $0=x(e)<y(e)$. In both cases, we obtain $e\in E^{>}$, completing the examination of ((4.22)).
Now considering the relations in ((4.21)) and ((4.22)), we can conclude that $C_{w}(y_{w}\vee x^{\prime}_{w})=y_{w}$. Thus, $y_{w}\succeq_{w}x^{\prime}_{w}$, and this relation holds for all $w\in V^{\gtrless}\cap W$, yielding $y\succeq_{W}x^{\prime}$. By the polarity, this implies $y\preceq_{F}x^{\prime}$, proving the lemma. $\Box$
[…]
In fact, the above analysis enables us to establish additional useful properties. First of all the following sharper version of Lemma 4.11 can be extracted from the above proof (it will be used in Sect 5).
Lemma 4.12
Let $x,y\in{\cal S}$ and $y\prec x$. Then there is a maximal strong component $L$ in $\overrightarrow{\Gamma}(x)$ and a number $\lambda\leq\tau_{L}$ such that the assignment $x^{\prime}:=x+\lambda\rho_{L}$ satisfies $y\preceq x^{\prime}\prec x$. The maximum possible weight $\lambda$ (subject to $y\preceq x^{\prime}$) is equal to the minimum between the number $\tau_{L}$, the values $(y(e)-x(e))/\rho_{L}(e)$ over $e\in E_{L}$ such that $\rho_{L}(e)>0$ and $y(e)>x(e)$, and the values $(x(e)-y(e))/|\rho_{L}(e)|$ over $e\in E_{L}$ such that $\rho_{L}(e)<0$ and $x(e)>y(e)$. normal-$\Box$
Also, using Lemma 4.11, one can conclude with the following
Corollary 4.13
For $x\in{\cal S}$, the following are equivalent:
  • (i)
    $x\neq x^{\rm max}$;
  • (ii)
    for $\varepsilon>0$, system (4.5)–(4.8) has a solution;
  • (iii)
    the set of active edges (see Definition 4) is nonempty;
  • (iv)
    the active graph $\overrightarrow{\Gamma}(x)$ has a strongly connected component.
Indeed, if (i) is valid, then for $y:=x^{\rm max}$, we have $y\prec x$, and Lemma 4.11 implies validity of (ii). Conversely, if system (4.5)–(4.8) is solvable for $\varepsilon>0$, then (by Lemma 4.8) there exists a stable $x^{\prime}$ such that $x^{\prime}\prec x$, yielding (i). The equivalence of (iii) and (iv) can be concluded from Cleaning procedure. The implication (iii)$\to$(ii) follows from Lemma 4.4. The converse implication (ii)$\to$(iii) can be concluded from the discussion in the beginning part of this section.
5 Additional properties of rotations
In this section we discuss additional (in particular, computational) aspects related to rotations and expose an appealing example of rotations.
We will refer to the transformation $x\rightsquigarrow x^{\prime}:=x+\lambda\rho_{L}$, where $L\in{\cal L}^{\rm rot}(x)$ and $\lambda$ is an admissible (w.r.t. $x$) weight for the rotation $\rho_{L}$, as the rotational shift along $\rho_{L}$ of value $\lambda$. For brevity, we also say that $x^{\prime}$ is obtained from $x$ by the shift by $\lambda\rho_{L}$. If $\lambda=\tau_{L}(x)$, the shift is called full. As before, the sets $F\cap V_{L}$ and $W\cap V_{L}$ are denoted as $F_{L}$ and $W_{L}$, respectively. One can see that
  1. (5.1)
    the maximal admissible weight $\tau_{L}=\tau_{L}(x)$ is equal to the minimum of the following values: (a) $x(e)/|\rho_{L}(e)|$ over all $e\in H_{w}$, $w\in W_{L}$; (b) $(b(e)-x(e))/\rho_{L}(e)$ over all $e\in D_{f}$, $f\in F_{L}$; and (c) $(x(e)-x(e^{\prime}))/(|\rho_{L}(e)|+\rho_{L}(e^{\prime}))$ over all pairs $e,e^{\prime}\in E_{L}$ such that $e\in H_{w}$ and $e^{\prime}\in\pi^{\rm c}_{w}(x)-H_{w}$ among $w\in W_{L}$.
(Here (a) and (b) ensure that $x^{\prime}$ is nonnegative and does not exceed the upper bounds $b$, and (c) provides (when the critical tie preserves) the inclusion $H_{w}(x)\subseteq H_{w}(x^{\prime})$ for $w\in W_{L}$. In other words, in case (c), the increased assignment in any edge of the set $\pi^{\rm c}_{w}-H_{w}$ should not “overlap” the decreased assignments in $H_{w}$.)
Given a stable $x$, the extraction of critical ties and their heads, and, further, the construction of the active graph $\overrightarrow{\Gamma}$, take linear time $O(|E|)$ (assuming that $|V|=O(|E|)$). The task of finding the strong components in $\overrightarrow{\Gamma}$ can also be solved in linear time (the first linear time algorithm for this task is given in [11]). Then we can easily extract the set ${\cal L}^{\rm rot}$ of maximal strong components. For each $L\in{\cal L}^{\rm rot}$, we can efficiently construct the rotation $\rho_{L}$ by applying to system (4.5)–(4.7) a strongly polynomial method in linear algebra (see, e.g., [15]). (At this moment we cannot suggest a more combinatorial efficient method to solve this system.) Finally, using ((5.1)), it is not difficult to compute the maximal rotational weights $\tau_{L}$ for all rotations $\rho_{L}$ in time $O(|E|)$ (taking into account that the sets $V_{L}$ are pairwise disjoint).
As a result, we obtain the following
Proposition 5.1
Given $x\in{\cal S}$, we can perform, in time polynomial in $|E|$, the following tasks: (a) recognize whether $x=x^{\rm max}$ or not, and (b) in case $x\neq x^{\rm max}$, find all rotations $\rho_{L}$, $L\in{\cal L}^{\rm rot}(x)$, and their maximal admissible weights $\tau_{L}$.
Next, based on the analysis from the previous section, one can estimate the number of rotational transformations (shifts) applied in the process of moving from the minimal stable assignment $x^{\rm min}$ to an arbitrary stable $x\in{\cal S}$, in particular, to the maximal assignment $x^{\rm max}$. Here the following observations are of use (cf. ((4.15)) and ((5.1))):
  1. (5.2)
    Let $x^{\prime}$ be obtained from $x\in{\cal S}$ by the shift along the rotation $\rho_{L}$ of value $\lambda$; then:
    • (a)
      for each $f\in F^{=}$, either $\pi^{\rm c}_{f}(x^{\prime})=\pi^{\rm c}_{f}(x)$ and $H_{f}(x^{\prime})\subseteq H_{f}(x)$, or $\pi^{\rm c}_{f}(x^{\prime})$ is worse than $\pi^{\rm c}_{f}(x)$; in its turn, for each $w\in W^{=}$, either $\pi^{\rm c}_{f}(x^{\prime})=\pi^{\rm c}_{f}(x)$ and $H_{w}(x^{\prime})\supseteq H_{w}(x)$, or $\pi^{\rm c}_{w}(x^{\prime})$ is better than $\pi^{\rm c}_{w}(x)$;
    • (b)
      if $\lambda=\tau_{L}$, then at least one of the following takes place: (i) for some $f\in F^{=}$, either $\pi^{\rm c}_{f}(x^{\prime})=\pi^{\rm c}_{f}(x)$ and $H_{f}(x^{\prime})\subset H_{f}(x)$, or $\pi^{\rm c}_{f}(x^{\prime})<\pi^{\rm c}_{f}(x)$; (ii) for some $w\in W^{=}$, either $\pi^{\rm c}_{w}(x^{\prime})=\pi^{\rm c}_{w}(x)$ and $H_{w}(x^{\prime})\supset H_{w}(x)$, or $\pi^{\rm c}_{w}(x^{\prime})>\pi^{\rm c}_{w}(x)$.
This monotonicity behavior leads to the following:
  1. (5.3)
    in the process of rotational shifts, for each edge $e\in fw\in E$, the updates of the current assignment $x(e)$ can be chronologically divided into two phases: at the first phase, the value at $e$ is monotone non-decreasing (when $e$ occurs in “heads” concerning the part $F$); and at the second phase, the value at $e$ is monotone non-increasing (when $e$ occurs in “heads” concerning the part $W$).
Using ((5.2)) and ((5.3)), we can obtain the following
Lemma 5.2
(i) If $x\in{\cal S}$ is constructed from $x^{\rm min}$ by a sequence of full rotational shifts, then the number of shifts is $O(|E|)$.
(ii) Any $x\in{\cal S}$ can be constructed from $x^{\rm min}$ by $O(|E|)$ rotational shifts.
Proof  To show (i), let us associate with a full shift in a given sequence one edge $e=fw$ along with some of the following events: 1) $e$ is inserted initially in the head $H_{f}$, where $f\in F$; 2) $e$ leaves the head $H_{f}$; 3) $e$ is inserted initially in the head $H_{w}$, where $w\in W$; and 4) $e$ leaves the head $H_{w}$. More precisely, if $x^{\prime}$ is obtained from $x$ by a full shift along a rotation $\rho_{L}$, then, considering the corresponding cases in ((5.2))(b), we choose $e\in H_{f}(x)-H_{f}(x^{\prime})$ in case $H_{f}(x^{\prime})\subset H_{f}(x)$; choose $e\in\pi^{\rm c}_{f}(x^{\prime})$ in case $\pi^{\rm c}_{f}(x^{\prime})<\pi^{\rm c}_{f}(x)$; choose $e\in H_{w}(x^{\prime})-H_{w}(x)$ in case $H_{w}(x^{\prime})\supset H_{w}(x)$; and choose $e\in H_{w}(x)$ in case $\pi^{\rm c}_{w}(x^{\prime})>\pi^{\rm c}_{w}(x)$. Then these choices match the events 1)–4), and one can see that each edge is associated with at most four full shifts in the process.
To show (ii), we first notice that any stable $x\in{\cal S}$ can be obtained from $x^{\rm min}$ by a finite sequence of rotational shifts with some admissible weights (this can be seen, e.g., from Lemma 4.12). If in this sequence, one and the same rotation $\rho_{L}$ (or, better to say, the same strong component $L$ in current active graphs) occurs several times, then we can combine the shifts related to these occurrences into one shift along $\rho_{L}$, summing up the corresponding rotational weights. (Here we use the fact that if a shift along $\rho_{L}$ is not full, then $L$ continues to be a maximal strong component in the current active graph, cf. the proof of Proposition 4.9.) Thus, the initial sequence can be transformed into another one, going from $x^{\rm min}$ to $x$ as before, so that each rotation in it occur exactly once. Then the rotations with non-maximal weights are related to different maximal strong components in the last graph $\overrightarrow{\Gamma}(x)$), and therefore, the number of these is $O(|V|)$. Now the desired estimate follows from (i). $\Box$
Remark 2. In fact, a slightly more refined analysis shows that the term $O(|E|)$ figured (twice) in Lemma 5.2 can be replaced by $2(|E|)$ (we omit details here). A similar result (with the bound $2|E|$) for SAP was established in [4] (extending an earlier, well known, result of this sort for the stable marriage problem in [9]).
Example 3. As is mentioned in Remark 1 from Sect. 4, there exist rotations with “big” values. An example of this sort is shown in Fig. 1. Here we demonstrate an active graph $\Gamma$ with vertex parts $F=\{f_{1},f_{2},f_{3}\}$ and $W=\{w_{1},w_{2},w_{3}\}$, the edges active for $F$ ($W$) are drawn by bold (resp. dotted) lines, and the numbers on edges point out values of rotation $\rho$ (this is a normalized aligned $Z$-circulation); they vary in the interval $[-8,7]$.
This rotation appears in the following instance of SMP: $G=\Gamma$; the weak orders at vertices of $G$ are defined as:
for vertex $f_{i}$: $f_{i}w_{1}\sim f_{i}w_{2}\sim f_{i}w_{3},\quad i=1,2,3;$
for vertex $w_{1}$: $f_{1}w_{1}>f_{3}w_{1}>f_{2}w_{1};$
for vertex $w_{2}$: $f_{1}w_{2}>f_{2}w_{2}>f_{3}w_{2};$
for vertex $w_{3}$: $f_{2}w_{3}>f_{1}w_{3}\sim f_{3}w_{3}.$
For arbitrarily chosen numbers $0<a<b$, the upper bound $b(e)$ for each edge $e$ is equal to $b$, and the quota $q(v)$ of each vertex $v$ is equal to $3a$. Let $x(e):=a$ for all edges $e$ of $G$. One can see that: the assignment $x$ is fully filled for all vertices ($F^{=}(x)=F$ and $W^{=}(x)=W$); for all $w_{j}\in W$, the critical ties $\pi^{\rm c}_{w_{j}}(x)$ consist just of the dotted edges and they coincide with the heads $H_{w_{j}}(x)$; for all $f_{i}\in F$, the reduced heads $\widetilde{H}_{f_{i}}(x)$ consist of the bold edges. As a result, $x$ is stable. All edges are non-saturated by $x$, and one can see that the directed active graph $\overrightarrow{\Gamma}(x)$ has a unique strong component $L$ involving all edges. Then $\rho_{L}=\rho$. Applying ((5.1)), we obtain that the maximum rotational weight $\tau_{L}$ is equal to the minimum of $a/8$ and $(b-a)/7$.
Figure 1: A rotation with “big” values
6 The poset of rotations
In this section we explain that the stable assignments in our mixed model admit an efficient representation, via the so-called closed sets (weighted analogs of ideals) in a poset generated by the rotations.
We have seen that any stable assignment $x\in{\cal S}$ can be obtained from $x^{\rm min}$ (which is the best for $F$ and the worst for $W$) by use of a sequence of $O(|E|)$ rotational shifts, see Lemma 5.2. The case $x=x^{\rm max}$ is of an especial interest for us.
Definition 7. Let $T$ be a sequence $x_{0},x_{1},\ldots,x_{N}$ of stable assignments such that each $x_{i}$ is obtained from $x_{i-1}$ by a rotational shift, i.e., $x_{i}=x_{i-1}+\lambda_{i}\rho_{L_{i}}$ (where $\rho_{L_{i}}$ is the rotation generated by a element (maximal strong component) $L_{i}$ of the set ${\cal L}^{\rm rot}(x_{i-1})$, and $\lambda_{i}$ is an admissible weight: $\lambda_{i}\leq\tau_{L_{i}}(x_{i-1})$; see Definition 6. We say that $T$ is non-expensive is all rotations $\rho_{L_{i}}$ are different (as vectors in ${\mathbb{R}}^{E}$). In particular, $T$ is non-expensive if all shifts involved are full, i.e., each $\lambda_{i}$ is the maximal rotational weight for $L_{i}$ (at the moment it applies). This always happens (in a non-expensive case) when $x_{0}=x^{\rm min}$ and $x_{N}=x^{\rm max}$; in this case we call $T$ a (complete) route. The set of all routs is denoted by ${\cal T}^{0}$.
When $x_{0}=x^{\rm min}$ and $x=x_{N}$ is arbitrary (resp. $x=x_{0}$ is arbitrary and $x_{N}=x^{\rm max}$), we call $T$ a lower (resp. upper) semi-route. The sequence of lower (resp. upper) sub-routes for $x\in{\cal S}$ is denoted by ${\cal T}^{\rm low}(x)$ (resp. ${\cal T}^{\rm up}(x)$). The set of rotations used to obtain a route or semi-route $T$ is denoted as ${\cal R}(T)$, and the set of pairs $(\lambda,\rho)$ formed by rotations $\rho\in{\cal R}(T)$ and their corresponding weights related to $T$ is denoted by $\Omega(T)$.
An important fact is that all routes use the same set of rotations and the same set of their weights. In other words, a strong component $L$ determining the rotation vector $\rho_{L}$ which is used in one route should necessarily appear in any other one and, moreover, for each rotation, the value of shift along it is always the same. Initially the phenomenon of this sort was revealed for rotations in the stable marriage problem in [9]; subsequently, this was established for other stability problems, in particular, for SAP in [4].
Generalizing a proving method from [4], we show a similar property for SMP.
Proposition 6.1
The set of pairs $\Omega(T)$ is the same for all routes $T\in{\cal T}^{0}$ (going from $x^{\rm min}$ to $x^{\rm max}$).
Proof  Let us say that a stable assignment $x\in{\cal S}$ is good if the set $\Omega(T)$ is the same for any upper semi-routes $T\in{\cal T}^{\rm up}(x)$; otherwise $x$ is called bad. We have to show that $x^{\rm min}$ is good.
Suppose that this is not so, and consider a maximal bad assignment $x$, in the sense that any $x^{\prime}$ obtained from $x$ by a rotational shift of maximal admissible weight is already good. Since $x$ is bad, in the set ${\cal T}^{\rm up}(x)$, there are two semi-routes $T^{\prime}=(x,x^{\prime},\ldots,x^{\rm max})$ and $T^{\prime\prime}=(x,x^{\prime\prime},\ldots,x^{\rm max})$ such that $\Omega(T^{\prime})\neq\Omega(T^{\prime\prime})$.
Let $x^{\prime}$ and $x^{\prime\prime}$ (occurring in $T^{\prime},T^{\prime\prime}$) be obtained from $x$ by rotational shifts by $\lambda^{\prime}\rho_{L^{\prime}}$ and $\lambda^{\prime\prime}\rho_{L^{\prime\prime}}$, respectively. These shifts involve different maximal strong components $L^{\prime}$ and $L^{\prime\prime}$ which are simultaneously contained in the active graph $\overrightarrow{\Gamma}(x)$. These components are disjoint, and therefore, the above rotational shifts commute. Then we can form in ${\cal T}(x)$ two semi-routes; namely, $\widetilde{T}^{\prime}$ starting with the triple $x,x^{\prime},x^{\prime}+\lambda^{\prime\prime}\rho_{L^{\prime\prime}}$, and $\widetilde{T}^{\prime\prime}$ starting with the triple $x,x^{\prime\prime},x^{\prime\prime}+\lambda^{\prime}\rho_{L^{\prime}}$, which further continue by the same upper semi-route $\widetilde{T}$ starting with $x+\lambda^{\prime}\rho_{L^{\prime}}+\lambda^{\prime\prime}\rho_{L^{\prime\prime}}$.
Now since the assignment $x^{\prime}$ is good (in view of the maximal choice of $x$), the parts of the semi-routes $T^{\prime}$ and $\widetilde{T}^{\prime}$ going from $x^{\prime}$ to $x^{\rm max}$ must determine the same sets of pairs $\Omega$, and similarly for the parts of $T^{\prime\prime}$ and $\widetilde{T}^{\prime\prime}$ going from $x^{\prime\prime}$. But then $\Omega(T^{\prime})=\{(\lambda^{\prime},\rho_{L^{\prime}}),\,(\lambda^{\prime\prime},\rho_{L^{\prime\prime}})\}\cup\Omega(\widetilde{T})=\Omega(T^{\prime\prime})$, contrary to the choice of $T^{\prime},T^{\prime\prime}$. $\Box$
Note that from this proposition we can easily obtain the following
Corollary 6.2
For any $x,y\in{\cal S}$ with $x\succ y$, the sets of pairs $(\lambda,\rho)$ used to form non-expensive sequences (“middle semi-routes”) from $x$ to $y$ are the same.
[…]
Definition 8.  A function $\lambda:{\cal R}\to{\mathbb{R}}_{+}$ is called closed if $\lambda(\rho)\leq\tau(\rho)$ for all $\rho\in{\cal R}$, and, in addition, $\lambda(\rho)>0$ and $\rho^{\prime}\lessdot\rho$ imply $\lambda(\rho^{\prime})=\tau(\rho^{\prime})$. A closed function $\lambda$ is called fully closed if for each $\rho\in{\cal R}$, the value $\lambda(\rho)$ is either 0 or $\tau(\rho)$. The set of closed functions for $({\cal R},\tau,\lessdot)$ is denoted by $\mathfrak{C}$. It forms a distributive lattice in which for closed $\lambda,\lambda^{\prime}$, the meet $\lambda\wedge\lambda^{\prime}$ and join $\lambda\vee\lambda^{\prime}$ are naturally defined by taking the coordinate-wise minimum and maximum, respectively.
Since the maximal weights $\tau(\rho)$ for all $\rho\in{\cal R}$ are strictly positive, we observe that for a closed function $\lambda$, both sets $I^{+}(\lambda):=\{\rho\colon\lambda(\rho)>0\}$ and $I^{-}(\lambda):=\{\rho\colon\lambda(\rho)=\tau(\rho)\}$ form ideals of the poset $({\cal R},\lessdot)$ (i.e. $I=I^{+}(\lambda)$ is such that $\rho\in I$ and $\rho^{\prime}\lessdot\rho$ imply $\rho^{\prime}\in I$, and similarly for $I^{-}(\lambda)$).
A relationship between the stable assignments and closed functions is established as follows. For $x\in{\cal S}$, consider a lower semi-route $T=(x^{\rm min}=x_{0},x_{1},\ldots,x_{N}=x)$, and for $i=1,\ldots,N$, let $x_{i}$ be obtained from $x_{i-1}$ by a rotational shift $\lambda_{i}\rho_{i}$. By Corollary 6.2, the set $\Omega(T)=\{(\lambda_{i},\rho_{i})\colon i=1,\ldots,N)\}$ does not depend on a semi-route in ${\cal T}^{\rm low}(x)$. Moreover, the function $\lambda:{\cal R}\to{\mathbb{R}}_{+}$, where $\lambda(\rho)$ is defined to be $\lambda_{i}$ if $\rho=\rho_{i}$, $i=1,\ldots,N$, and 0 otherwise, is closed. (Indeed, if $\rho=\rho_{i}$ and $\rho^{\prime}\lessdot\rho$, then $\rho^{\prime}$ must be used in any semi-route in ${\cal T}^{\rm low}(x)$, and its generating strong component $L^{\prime}$ cannot be in ${\cal L}^{\rm rot}(x_{i})$ (for otherwise $\rho^{\prime}$ and $\rho$ can be applied in any order). It follows that $\lambda(\rho^{\prime})=\tau(\lambda^{\prime})$, as required.) We define this closed function by $\omega(x)$.
Proposition 6.3
The correspondence $x\mapsto\omega(x)$ determines a bijection between the lattice $({\cal S},\prec_{W})$ of stable assignments and the lattice $(\mathfrak{C},<)$ of closed functions for $({\cal R},\tau,\lessdot)$.
Proof Consider stable assignments $x,y\in{\cal S}$. Let $\lambda:=\omega(x)$ and $\mu:=\omega(y)$. The key part of the proof is to show that
$\lambda\wedge\mu=\omega(x\curlywedge y)\qquad\mbox{and}\qquad\lambda\vee\mu=\omega(x\curlyvee y)$ (6.1)
[…]
To show this, consider a middle (non-expensive) semi-route $X=(z=x_{0},x_{1},\ldots,x_{N})$ from $z:=x\curlywedge y$ to $x$, and a middle semi-route $Y=(z=y_{0},y_{1},\ldots,y_{M})$ from $z$ to $y$. Let ${\cal R}(X)=(\rho_{1},\ldots,\rho_{N})$ and ${\cal R}(Y)=(\rho^{\prime}_{1},\ldots,\rho^{\prime}_{M})$ be the corresponding sequences of rotations, and let $\lambda_{i}$ ($\lambda^{\prime}_{j}$) be the corresponding weight of $\rho_{i}$ (resp. $\rho^{\prime}_{j}$) used in $X$ (resp. $Y$). (By Corollary 6.2, the set of rotations $\rho_{i}$ in ${\cal R}(X)$, along with their weights $\lambda_{i}$, depend only on the end assignments $z$ and $x$. And similarly for ${\cal R}(Y)$.)
One can realize that the first relation in (6.1) follows from the property that
  1. (6.2)
    the sequences ${\cal R}(X)$ and ${\cal R}(Y)$ have no common rotations.
We prove ((6.2)) as follows. Suppose, for a contradiction, that ${\cal R}(X)$ and ${\cal R}(Y)$ share a rotation $\rho$, say, $\rho=\rho_{i}=\rho^{\prime}_{j}$.
If $i=j=1$ (i.e., $\rho$ is the first rotation in both ${\cal R}(X)$ and ${\cal R}(Y)$), then the assignment $z^{\prime}:=z+\min\{\lambda_{1},\lambda^{\prime}_{1}\}\,\rho$ satisfies $z^{\prime}\succ z$, $z^{\prime}\preceq x$ and $z^{\prime}\preceq y$, contrary to the fact that $z$ is the maximal lower bound for $x,y$. So we may assume that $i>1$ and $j\geq 1$.
Consider a rotation $\rho_{k}$ with $k<i$ in ${\cal R}(X)$. If $\rho_{k}\lessdot\rho$, then $\rho_{k}$ must occur in the sequence ${\cal R}(Y)$ earlier than $\rho^{\prime}_{j}=\rho$, and therefore, $\rho_{k}=\rho^{\prime}_{k^{\prime}}$ for some $k^{\prime}<j$. This enables us to exclude the situation when
  • ($\ast$):
    $\rho_{k}\lessdot\rho_{i}$ for $k=1,\ldots,i-1$, and simultaneously, $\rho^{\prime}_{k^{\prime}}\lessdot\rho^{\prime}_{j}$ for $k^{\prime}=1,\ldots,j-1$.
Indeed, in this case, we would have $\{\rho_{1},\ldots,\rho_{i}\}=\{\rho^{\prime}_{1},\ldots,\rho^{\prime}_{j}\}$. Also for each $k=1,\ldots,i-1$, the relation $\rho_{k}\lessdot\rho$ implies that $\lambda_{k}$ is the maximal admissible weight $\tau(\rho_{k})$. It follows that if $\rho_{k}=\rho^{\prime}_{k^{\prime}}$, then $\lambda_{k}=\lambda^{\prime}_{k^{\prime}}$. Then the stable assignment $z^{\prime}:=z+\lambda_{1}\rho_{1}+\cdots+\lambda_{i-1}\rho_{i-1}$ belongs to both semi-routes $X$ and $Y$. This implies $z\prec z^{\prime}\preceq x$ and $z^{\prime}\preceq y$, contradicting the definition of $z$.
It remains to consider the situation when there is $\rho_{k}$ with $1\leq k<i$ which is not comparable with $\rho_{i}$; for brevity, we will write $\rho_{i}\approx\rho_{k}$. We will use the following
Claim. Let $\widetilde{x}\in{\cal S}$ and let $\rho$ and $\rho^{\prime}$ be rotations such that $\rho$ is applicable to $\widetilde{x}$, and $\rho^{\prime}$ is applicable to $\widetilde{x}^{\prime}:=\widetilde{x}+\lambda\rho$ for an admissible weight $\lambda$ for $\rho$. Then exactly one of the following is valid: (a) $\rho^{\prime}$ is applicable to $\widetilde{x}$, or (b) $\rho\lessdot\rho^{\prime}$.
This Claim will be proved later. Assuming validity of it, we argue as follows. Let $k$ be maximum subject to $k<i$ and $\rho_{k}\approx\rho_{i}$. First suppose that $k=i-1$. Then, by (a) in the Claim applied to $\widetilde{x}=x_{i-2}$, the rotations $\rho^{\prime}=\rho_{i-1}$ and $\rho=\rho_{i}$ commute, and we can update $X$ by swapping $\rho^{\prime}$ and $\rho$. This gives the semi-route $X^{\prime}=(x_{0},\ldots,x_{i-2},x^{\prime},x_{i},\ldots,x_{N})$, where $x^{\prime}:=x_{i-2}+\lambda_{i}\rho_{i}$, and in the corresponding sequence ${\cal R}(X^{\prime})$ the rotation $\rho$ is disposed earlier than in ${\cal R}(X)$. We can continue to further “lower” $\rho$ if $\rho_{i-2}\approx\rho$, and so on.
Now suppose that $k<i-1$. Then we are going to “lift” $\rho^{\prime}=\rho_{k}$ so as to get closer to $\rho$. This is done as follows. Since $\rho^{\prime}\approx\rho$ and $\rho_{k+1}\lessdot\rho$ (by the maximal choice of $k$), we have $\rho^{\prime}\approx\rho_{k+1}$ (for $\rho^{\prime}\lessdot\rho_{k+1}$ would imply $\rho^{\prime}\lessdot\rho$). Then, by (a) in the Claim, $\rho_{k+1}$ is applicable to $x_{k-1}$. Swapping $\rho^{\prime}$ and $\rho_{k+1}$ (preserving their weights), we obtain the semi-route $X^{\prime\prime}$ such that in the corresponding sequence ${\cal R}(X^{\prime\prime})$ the rotation $\rho^{\prime}$ becomes closer to $\rho$, as required.
Summing up the above observations, we can conclude that the semi-route $X$ can be transformed so that the rotation $\rho$ reaches the first position in the resulting sequence of rotations related to $z$ and $x$. Transforming $Y$ in a similar way, we obtain a sequence of rotations related to $z$ and $y$ in which the same rotation $\rho$ is situated at the first position as well. Then we act as in the beginning of the proof, yielding a contradiction with the definition of $z$. This proves ((6.2)), implying the left relation in (6.1). The right relation in (6.1) is proved symmetrically (by taking the order $\prec_{F}$ instead of $\prec_{W}$, and thinking of a route as a sequence going from $x^{\rm max}$ to $x^{\rm min}$). Thus, $\omega$ gives a homomorphism of the above lattices as required.
Now to see that $\mathfrak{C}$ is bijective to ${\cal S}$, it suffices to note that for any two rotations $\rho$ and $\rho^{\prime}$, there is $x\in{\cal S}$ which “separates” these rotations, in the sense that a lower semi-route for $x$ uses one of $\rho,\rho^{\prime}$, while an upper semi-route for $x$ uses the other of them. $\Box$
Proof of the Claim. Let $L$ and $L^{\prime}$ be the maximal strong components in the active graphs $\overrightarrow{\Gamma}=\overrightarrow{\Gamma}(\widetilde{x})$ and $\overrightarrow{\Gamma}^{\prime}=\overrightarrow{\Gamma}(\widetilde{x}^{\prime})$ that determine $\rho$ and $\rho^{\prime}$, respectively. If $\rho^{\prime}$ is applicable to $\widetilde{x}$, or, equivalently, $L^{\prime}\in{\cal L}^{\rm rot}(\widetilde{x})$, then $\rho\approx\rho^{\prime}$. So assume that $L^{\prime}\notin{\cal L}^{\rm rot}(\widetilde{x})$. Then $\lambda$ is the maximal rotational weight $\tau(\rho)$, whence $\overrightarrow{\Gamma}\neq\overrightarrow{\Gamma}^{\prime}$.
For $v\in V$, let $A_{v}$ and $A^{\prime}_{v}$ denote the sets of active edges leaving $v$ in $\overrightarrow{\Gamma}$ and $\overrightarrow{\Gamma}^{\prime}$, respectively. Then (cf. ((5.2))(b)):
  1. (6.3)
    if $v=f\in F$, then either $\pi^{\rm c}_{f}(\widetilde{x}^{\prime})=\pi^{\rm c}_{f}(\widetilde{x})$ and $A^{\prime}_{f}\subseteq A_{f}$, or $\pi^{\rm c}_{f}(\widetilde{x}^{\prime})$ is worse than $\pi^{\rm c}_{f}(\widetilde{x})$; whereas if $v=w\in W$, then either $\pi^{\rm c}_{w}(\widetilde{x}^{\prime})=\pi^{\rm c}_{w}(\widetilde{x})$ and $A^{\prime}_{w}\supseteq A_{w}$, or $\pi^{\rm c}_{w}(\widetilde{x}^{\prime})$ is better than $\pi^{\rm c}_{w}(\widetilde{x})$; moreover, if $\pi^{\rm c}_{v}(\widetilde{x}^{\prime})=\pi^{\rm c}_{v}(\widetilde{x})$ for all $v\in V$, then at least one of the above inclusions is strict.
Now consider two possible cases.
Case 1: $L$ and $L^{\prime}$ have a common vertex. For each $v\in V_{L}\cap V_{L^{\prime}}$, we have: $A_{v}=E_{L}\cap E_{v}$ and $A^{\prime}_{v}=E_{L^{\prime}}\cap E_{v}$, the sets $A_{v},\,A^{\prime}_{v},\,\pi^{\rm c}_{v}(\widetilde{x}),\,\pi^{\rm c}_{v}(\widetilde{x}^{\prime})$ satisfy relations as in ((6.3)), and at least one of these relations is different from equality (since, obviously, $L\neq L^{\prime}$). When moving along any route, relations concerning the active graphs for neighboring assignments are updated similar to those exposed in ((6.3)). These updates have a “monotone character” at each vertex, and we can conclude that the rotation $\rho^{\prime}$ can never be applied in a route earlier than $\rho$. It follows that $\rho\lessdot\rho^{\prime}$.
Case 2: $L$ and $L^{\prime}$ are disjoint. This is possible only if $L^{\prime}$ is a strong component in each of $\overrightarrow{\Gamma}$ and $\overrightarrow{\Gamma}^{\prime}$, but $L^{\prime}$ is maximal in $\overrightarrow{\Gamma}^{\prime}$, and not maximal in $\overrightarrow{\Gamma}$. Then the set $Q$ of active edges in $\overrightarrow{\Gamma}$ going from $L^{\prime}$ to other strong components in $\overrightarrow{\Gamma}$ must disappear under the transformation of $\widetilde{x}$ along $\rho$. This implies that all edges in $Q$ enter just $L$. Since the assignment on each $e\in Q$ is not changed: $\widetilde{x}^{\prime}(e)=\widetilde{x}(e)$, the situation that the edge $e$ becomes non-active can happen only if $e$ goes from $F$ to $W$, say, $e=(f,w)$, and $w$ does not belong to the tail $T_{w}(\widetilde{x}^{\prime})$. This implies that the assignments on the edges in $\pi^{\rm c}_{w}(\widetilde{x})$ must reduce to zero; then the critical tie in $\Pi_{w}$ becomes better: $\pi^{\rm c}_{w}(\widetilde{x}^{\prime})>\pi^{\rm c}_{w}(\widetilde{x})$. Moreover, $e$ must belong to a tie worse than $\pi^{\rm c}_{w}(\widetilde{x}^{\prime})$. It follows that in any route the rotation $\rho^{\prime}$ cannot be applied earlier than $\rho$ (for otherwise the edges from $Q$ should be active at the moment of applying $\rho^{\prime}$, and therefore, $L^{\prime}$ could not be a maximal strong component in the current active graph). As a result, we obtain $\rho\lessdot\rho^{\prime}$. $\Box$
In the rest of this section we explain how to construct the rotation poset $({\cal R},\tau,\lessdot)$ efficiently. Our method is of a similar spirit as that elaborated for the stable allocation problem in [4] (and earlier for the stable marriage problem in [9]).
First of all we compute, step by step, the set ${\cal R}$ of rotations, see explanations around Proposition 5.1. This is performed in parallel with the construction of a route $T$ from $x^{\rm min}$ to $x^{\rm max}$. Recall that at each step, for a current stable assignment $x$, we construct the active graph $\overrightarrow{\Gamma}(x)$, extract the set ${\cal L}^{\rm rot}(x)$ of maximal strong components $L$ in it, and determine the rotations $\rho=\rho_{L}$ and their maximal admissible weights $\tau(\rho)=\tau_{L}$. (Here finding ${\cal L}^{\rm rot}(x)$ can be performed in linear time $O(|E|)$, while computing the rotations $\rho_{L}$, which is a more involved task, takes time polynomial in $|E|$.) Then we update the current $x$ by adding to it the total shift $\sum(\tau_{L}\rho_{L}\colon L\in{\cal L}^{\rm rot}(x))$, and so on. As a result, we obtain the set ${\cal R}$ of all rotations $\rho$ (where $|{\cal R}|=O(|E|)$), along with their maximal weights $\tau(\rho)$, in time polynomial in $|E|$.
[…]
More precisely, define ${\cal E}$ to be the set of pairs $(\rho,\rho^{\prime})$ in ${\cal R}$ such that
  1. (6.4)
    there exists $x\in{\cal S}$ such that $\rho$ but not $\rho^{\prime}$ is applicable to $x$, whereas $\rho^{\prime}$ is applicable to the assignment $x^{\prime}:=x+\tau(\rho)\rho$.
One can see that each pair $(\rho,\rho^{\prime})$ in ${\cal E}$ gives an immediately preceding relation. (Indeed, $x$ and $x^{\prime}$ as above can be represented as two consecutive elements in some route; also the Claim implies that $\rho\lessdot\rho^{\prime}$; therefore, the existence of $\rho^{\prime\prime}$ between $\rho$ and $\rho^{\prime}$ is impossible.)
Note that the set of pairs ${\cal E}$ can be found efficiently (in fact, in time $O(|E|^{2})$).
We show that ${\cal E}$ does determine all immediately preceding relations in $({\cal R},\lessdot)$ (and therefore, the directed graph $({\cal R},{\cal E})$ can be used to encode this poset and, moreover, to store the information about all stable assignments). Clearly if $\rho^{\prime}\in{\cal R}$ is reachable from $\rho\in{\cal R}$ by a directed path in $({\cal R},{\cal E})$, then $\rho\lessdot\rho^{\prime}$. The converse is valid as well.
Proposition 6.4
For $\rho,\rho^{\prime}\in{\cal R}$, if $\rho\lessdot\rho^{\prime}$, then the graph ${\cal H}=({\cal R},{\cal E})$ has a directed path from $\rho$ to $\rho^{\prime}$.
Proof Let ${\cal I}$ be the set of ideals in the poset $({\cal R},\lessdot)$, and ${\cal J}$ the set of ideals $I$ in the poset determined by ${\cal H}$ (via the reachability by paths). Then ${\cal I}\subseteq{\cal J}$, and we have to show that ${\cal I}={\cal J}$, whence the result will follow.
Suppose this is not so. Choose $J\in{\cal J}$ with $|J|$ maximum subject to $J\notin{\cal I}$. Also choose $R\in{\cal I}$ with $|R|$ maximum subject to $R\subset J$. By Proposition 6.3, $R$ corresponds to some $x\in{\cal S}$, and therefore, $R$ is the set of rotations used to form a lower semi-route $T$ from $x^{\rm min}$ to $x$ (i.e., $R={\cal R}(T)$).
Consider a rotation $\rho$ applicable to $x$. Then $R^{\prime}:=R\cup\{\rho\}$ is an ideal in ${\cal I}$ (corresponding to the assignment $x^{\prime}:=x+\tau(\rho)\rho$). By the maximality of $R$, $\rho$ is outside $J$. We have $J^{\prime}:=J\cup\{\rho\}\in{\cal J}$, and the maximality of $J$ implies $J^{\prime}\in{\cal I}$.
Now since $J^{\prime}\supset R^{\prime}$ and both $J^{\prime},R^{\prime}$ are ideals in ${\cal I}$, the set $J^{\prime}-R^{\prime}$ contains a rotation $\rho^{\prime}$ such that the set $R^{\prime\prime}:=R^{\prime}\cup\{\rho^{\prime}\}$ is an ideal in ${\cal I}$. At the same time, $J^{\prime\prime}:=R\cup\{\rho^{\prime}\}$ is not an ideal in ${\cal I}$; for otherwise, in view of $R\subset J^{\prime\prime}\subseteq J$, we would obtain a contradiction with the maximality of $R$ (if $J^{\prime\prime}\neq J$) or with the condition on $J$ (if $J^{\prime\prime}=J$).
So we come to the situation when $R,R^{\prime},R^{\prime\prime}\in{\cal I}$, $R^{\prime}=R\cup\{\rho\}$, $R^{\prime\prime}=R^{\prime}\cup\{\rho^{\prime}\}$, but $R\cup\{\rho^{\prime}\}\notin{\cal I}$. Then the rotation $\rho^{\prime}$ is not applicable to the assignment $x$, but it is applicable to $x^{\prime}$ (obtained from $x$ by applying $\rho$). This means that $(\rho,\rho^{\prime})$ is an edge in ${\cal E}$ (cf. ((6.4))). But this edge goes from ${\cal R}-J$ to $J$ (since $\rho\notin J$ while $\rho^{\prime}\in J$), contrary to the fact $J$ is an ideal for ${\cal H}$. $\Box$
Appendix A Appendix. An efficient algorithm for finding $x^{\rm min}$
Recall that the method described in Sect. 3 can have infinitely many iterations, though it converges just to the stable assignment $x^{\rm min}$ which is optimal for the part $F$. We refer to that method, as well as iterations in it, as ordinary. In this section we describe a modified method, which is finite and finds $x^{\rm min}$ in time polynomial in $|E|$.
We start with analyzing some details of the ordinary method. Recall that a current, $i$-th, iteration in it consists of two stages: at the 1st one, the current boundary function $b^{i}$ is transformed into $x^{i}$, by setting $x^{i}_{f}:=C_{f}(b^{i}_{f})$ for each $f\in F$, and at the 2nd one, $x^{i}$ is transformed into $y^{i}$, by setting $y^{i}_{w}:=C_{w}(x^{i}_{w})$ for each $w\in W$. Then the boundary function is updated according to rule (3.1). Let $F^{i=}$ denote the set of fully filled vertices $f$ in $F$ for $x^{i}$ (i.e., $|x^{i}_{f}|=q(f)$), and $W^{i=}$ the set of fully filled vertices in $W$ for $y^{i}$. For $f\in F$, let $L^{i}_{f}$ denote the set of edges $e\in E_{f}$ such that $b^{i}(e)<b(e)$. We observe the following.
(i) For each $f\in F$, the sets $L_{f}$ are monotone non-decreasing: $L^{i}_{f}\subseteq L^{i+1}_{f}$, and the values on each edge $e\in L_{f}$ may decrease (at the 2nd stages of some iterations) but never increase (since $b^{1}(e)\geq\cdots\geq b^{i}(e)\geq b^{i+1}(e)\geq\cdots$).
(ii) The sets $W^{\bullet=}$ are monotone non-decreasing: $W^{i=}\subseteq W^{(i+1)=}$ (for if $|x^{i}_{w}|\geq q(w)$, then the choice function $C_{w}$ cuts $x^{i}_{w}$ so as to get $|y^{i}_{w}|=q(w)$, and subsequently the relations $|x_{w}|\geq|y_{w}|=q(w)$ are always maintained). In turn, the sets $F^{\bullet=}$ are monotone non-increasing: $F^{i=}\supseteq F^{(i+1)=}$ (since $b^{i}\geq b^{i+1}$).
(iii) The heads $H_{w}(y)$ of vertices $w\in W^{\bullet=}$ are monotone non-decreasing: $H_{w}(y^{i})\subseteq H_{w}(y^{i+1})$ (when $w\in W^{i=}$), unless the critical tie in $\Pi_{w}$ changes (becomes better). In turn, the heads $H_{j}(x)$ for $f\in F$ are monotone non-increasing: $H_{f}(x^{i})\supseteq H_{f}(x^{i+1})$ (taking into account that $H_{f}(x^{j})=\emptyset$ if $f\in F-F^{j=}$).
(iv) For each $f\in F$, all edges in $E_{f}-(H^{i}_{f}\cup L^{i}_{j})$ are saturated, where $H^{i}_{f}$ denotes the head $H_{f}(x^{i})$ (which is nonempty if $f\in F^{i=}$, and empty if $f\in F-F^{i=}$).
These observations point out monotone parameters, and we say that $i$-th iteration is positive if at least one of the following events happens in it: for some $f\in F$, the set of edges $e\in E_{f}$ that are either saturated or contained in $L_{f}$ increases; the set $W^{\bullet=}$ of fully filled vertices in $W$ increases; for some $w\in W^{\bullet=}$, either the head $H_{w}$ increases, or the critical tie for $w$ becomes better. The number of positive iterations does not exceed $2|E|+|W|$.
Next we deal with a sequence $Q$ of consecutive non-positive iterations and consider a current iteration among these. At the 1st stage of this iteration, the sum of values of the current function $x$ increases (compared with the preceding $y$) by some amount $\xi>0$. Note that if for a vertex $w\in W^{\bullet=}$, the value increases on some edge $e=fw$, then this edge cannot belong to the head $H_{w}$ or to a tie in $\Pi^{\rm end}_{w}$. (For otherwise the cutting operation at the 2nd stage of the current iteration decreases the value on $e$, and as a consequence, $e$ is added as a new element to the set $L_{f}$, which is impossible since the iteration is non-positive.)
It follows that an increase on some edge $e=fw$ at the 1st stage of an iteration in $Q$ is possible only in two cases: either $w\in W^{\bullet=}$ and $e$ is contained in the current tail $T_{w}$, or $w$ belongs to the current deficit set $W-W^{\bullet=}$. This implies that the total increase $\xi^{\prime}$ on the edges incident to vertices in $W^{\bullet=}$ satisfies $\xi^{\prime}\leq\xi$ (the equality may happen only if there is no contribution from $\xi$ to the vertices in $W-W^{\bullet=}$). At the 2nd stage of this iteration, the total value in the heads $H_{w}$ for these vertices $w\in W^{\bullet=}$ must decrease by the same amount $\xi^{\prime}$ (emphasize that these heads $H_{w}$ do not change, and each edge $fw$ where the value decreases has been added to the set $L_{f}$ on earlier iterations). Then we come to the 1st stage of the next iteration with the total “quota-underloading” $\xi^{\prime}$ in corresponding vertices in $F$, which further causes a “quota-overloading” $\xi^{\prime\prime}\leq\xi^{\prime}$ in $W^{\bullet=}$, and so on. When such “overloadings” decrease geometrically, non-positive iterations may continue infinitely long (yet providing a convergence to a limit function).
Relying on the observations above, we now are going to aggregate non-positive iterations from the ordinary method. More precisely, consider the first iteration in the sequence $Q$; assume that it is $i$-th iteration in the ordinary process. For brevity, we denote $x^{i}$ by $x$, and $y^{i}$ by $y$ (the functions obtained on the 1st and 2nd stages of the first iteration in $Q$). Also denote:
$F^{=}:=F^{i=}\quad\mbox{and}\quad W^{=}:=W^{i=};$
$H_{f}:=H_{f}(x),\;\;h_{f}:=|H_{f}|,\;\;\mbox{and}\;\;a(y,f):=y(e)\quad\mbox{for $f\in F^{=}$ and $e\in H_{f}$};$
$H_{w}:=H_{w}(y),\;\;h_{w}:=|H_{w}|,\;\;\mbox{and}\;\;a(y,w):=y(e)\quad\mbox{for $w\in W^{=}$ and $e\in H_{w}$}.$
Our idea consists in replacing the sequence $Q$ without the first iteration in it by one “big” (aggregated) iteration. To do so, we arrange the following linear program with variables $\phi_{f}$ for $f\in F^{=}$, and $\psi_{w}$ for $w\in W^{=}$:
maximize   $\eta(\phi):=\sum(h_{f}\phi_{f}\colon f\in F^{=})$    subject to (A.1)
$h_{w}\psi_{w}-\sum(\phi_{f}\colon f\in F^{=},\,fw\in H_{f})=0,\qquad w\in W^{=};$ (A.2)
$\qquad\qquad\qquad\psi_{w}\leq a(y,w),\qquad\qquad\qquad\qquad w\in W^{=};$ (A.3)
$h_{f}\phi_{f}-\sum(\psi_{w}\colon w\in W^{=},\,fw\in L_{f})\leq q(f)-|y_{f}|,\quad f\in F^{=};$ (A.4)
$\qquad\qquad\qquad\qquad\phi_{f}\leq b(fw)-y(fw),\qquad\qquad\qquad f\in F^{=},\;fw\in H_{f};$ (A.5)
$\qquad\phi_{f}+\psi_{w}\leq a(y,w)-y(fw),\qquad f\in F^{=},\,w\in W^{=},\,fw\in\pi^{\rm c}_{w}(y)-H_{w};$ (A.6)
$\qquad\sum(\phi_{f}\colon f\in F^{=},\,fw\in E_{w})\leq q(w)-|y_{w}|,\qquad\qquad w\in W-W^{=};$ (A.7)
$\phi\geq 0,\;\psi\geq 0.\qquad$ (A.8)
The set of feasible solutions to this system is nonempty (since we can take as $\phi,\psi$ the corresponding augmentations determined on the second iteration in $Q$) and compact (in view of (A.3),(A.5),(A.8)). Let $(\phi,\psi)$ be an optimal solution to (A.1)–(A.8), and let $\Delta=\Delta^{\phi,\psi}$ be the corresponding function on $E$, i.e., for $e=fw\in E$, $\Delta(e)$ is equal to $\phi_{f}$ if $f\in F^{=}$ and $e\in H_{f}$,  $-\psi_{w}$ if $w\in W^{=}$ and $e\in H_{w}$, and 0 otherwise.
Define $y^{\prime}:=y+\Delta^{\phi,\psi}$. Then, in light of the properties of $y$, we observe the following: (A.2) implies that all vertices in $W^{=}$ are fully filled for $y^{\prime}$; (A.3) implies that $y^{\prime}$ is nonnegative; (A.4) implies that $y^{\prime}$ does not violate the quotas within $F^{=}$; (A.5) implies that $y^{\prime}$ is does not exceed the upper bounds $b$; (A.6) can be rewritten as $y^{\prime}(e)=a(y,w)-\psi_{w}\geq y(fw)+\phi_{f}=y^{\prime}(fw)$, where $e\in H_{w}$, implying that the edges in the head $H_{w}$ for $y$ belong to the head at $w$ for $y^{\prime}$ as well; and (A.7) implies that $y^{\prime}$ does not violate the quotas within $W-W^{=}$.
One may assume that the optimal solution $(\phi,\psi)$ is chosen so that the set ${\cal L}$ of inequalities in (A.3)–(A.7) that turn into equalities is nonempty. One can see that if ${\cal L}$ involves a constraint from (A.3),(A.5)–(A.7), then some of the above-mentioned parameters changes. (Indeed, the equality with $w$ in (A.3) implies that $y^{\prime}$ becomes zero on $H_{w}(y)$, the equality with $fw$ in (A.5) implies that $y^{\prime}(fw)$ attains the upper bound $b(fw)$; the equality with $fw$ in (A.6) implies that the head at $w$ becomes larger; and the equality with $w$ in (A.7) implies that the vertex $w$ becomes fully filled.)
Now assume that none of the constraints in (A.3),(A.5)–(A.7) turns into equality, and therefore, the attained constraints concern only (A.4). Then the resulting function $y^{\prime}$ is well-defined, in the sense that it is suitable to applying further ordinary iterations. Let $x^{\prime\prime}$ be the function obtained by applying to $y^{\prime}$ the 1st stage of the ordinary iteration. We observe that if ${\cal L}$ involves all inequalities in (A.4), then all vertices in $F^{=}$ are fully filled for $y^{\prime}$; this implies that $x^{\prime\prime}=y^{\prime}$, and the process terminates.
Now suppose that some inequalities in (A.4) are strict for $y^{\prime}$, and therefore, the subset $F^{\prime}$ formed by $f\in F^{=}$ with $|y^{\prime}_{f}|<q(f)$ is nonempty. Let $y^{\prime\prime}$ be the functions obtained by applying to $x^{\prime\prime}$ the 2nd stage of the ordinary iteration. Then $\delta:=y^{\prime\prime}-y^{\prime}$ is represented as $\Delta^{\phi^{\prime},\psi^{\prime}}$, where $\phi^{\prime}$ is a function on $F^{=}$ such that $\phi^{\prime}_{f}>0$ for each $f\in F^{\prime}$. Now form the pair $(\phi^{\prime\prime},\psi^{\prime\prime})=(\phi+\phi^{\prime},\psi+\psi^{\prime})$. One can see that this $(\phi^{\prime\prime},\psi^{\prime\prime})$ is a feasible solution to (A.1)–(A.8). But $\eta(\phi^{\prime\prime})=\eta(\phi)+\eta(\phi^{\prime})>\eta(\phi)$, contrary to the optimality of $(\phi,\psi)$.
Thus, the big iteration either terminates the whole process or changes at least one of the above parameters. The modified method that we elaborate is organized so that each big iteration should be preceded by an ordinary iteration and there are no two consecutive ordinary non-positive iterations. Therefore, it finishes after $O(|E|)$ iterations. Each big iteration is performed in time polynomial in $|E|$, by using a strongly polynomial l.p. method (see, e.g., [15]) to solve linear program (A.1)–(A.8).
Let $\widehat{x}$ be the resulting assignment. To see that $\widehat{x}$ is stable, one can argue in a similar way as in the ordinary method (where a proof of stability is rather straightforward). More precisely, one can show by induction on the number of iterations that if an edge $e=fw$ is not saturated upon termination of a current iteration, then $e$ belongs to either the set $H_{f}\cup E^{\rm end}_{f}$ for $f\in F^{=}$, or to the set $H_{w}\cup E^{\rm end}_{w}$ for $w\in W^{=}$, whence the stability of $\widehat{x}$ will follow. (Here $W^{=},H_{w},E^{\rm end}_{w}$ are objects arising at the end of the iteration, whereas $F^{=},H_{f},E^{\rm end}_{f}$ are those arising on the 1st stage when the iteration is ordinary, and are defined as above (before introducing system (A.1)–(A.8)) when the iteration is big.)
It remains to ask: whether or not the resulting $\widehat{x}$ found by the modified algorithm does coincide with the assignment $x^{\rm min}$? In essence this question is not so important to us, because once we are given one or another stable assignment, we can always transform it into $x^{\rm min}$ by applying a sequence of $O(|E|)$ rotational shifts, by considering the polar order $\succeq_{W}$ on ${\cal S}$ and acting as in Sect. 6. As a result, we can conclude with the following
Proposition A.1
The modified algorithm is finite and finds the assignment $x^{\rm min}$. It consists of $O(|E|)$ iteration, each taking time polynomial in $|E|$. normal-$\Box$